### v4

# Shifted Contact Structures and Their Local Theory

Abstract
In this paper, we formally define the concept of $k$-shifted contact structures on derived (Artin) stacks and study their local properties in the context of derived algebraic geometry. In this regard, for negatively shifted contact derived $\mathbb{K}$-schemes, we develop a Darboux-like theorem and formulate the notion of symplectification.
1 Introduction and summary
Derived algebraic geometry (DAG) essentially provides a new setup to deal with non-generic situations in geometry (e.g. non-transversal intersections and “bad" quotients). To this end, it combines higher categorical objects and homotopy theory with many tools from homological algebra. Hence, roughly speaking, it can be considered as a higher categorical/homotopy theoretical refinement of classical algebraic geometry. In that respect, it offers a new way of organizing information for various purposes. Therefore, it has many interactions with other mathematical domains. For a survey of some directions, we refer to [1, 4].
In the context of DAG, it is also possible to work with familiar geometric structures, but in more general forms. For instance, $k$-shifted versions of Symplectic and Poisson geometries have already been described and studied in [6, 10]. In this regard, [7, 8, 9] offer some applications and local constructions.
Throughout this paper, we mainly work within the context of Toën & Vezzosi’s version of DAG [4, 5]. We also benefit from Lurie’s version [11]. In that respect, we always consider objects with higher structures in a functorial perspective, and we focus on nice representatives for those structures. For instance, by a derived ${\mathbb{K}}$-stack, we essentially mean a simplicial presheaf on the category of commutative differential graded $\mathbb{K}$-algebras (cdga) having nice local-to-global properties.
DAG provides an appropriate concept of a spectrum functor ${\rm Spec}$ from cdgas to (higher) spaces. Using this functor, we call a derived space of the form ${\bf X}\simeq{\rm Spec}A$ for some cdga $A$ an affine derived ${\mathbb{K}}$-scheme. As in the classical theory, a general derived ${\mathbb{K}}$-scheme $\bf Y$ is defined to be a space which is locally modeled on ${\bf X}\simeq{\rm Spec}A$. Note that affine derived schemes are in fact the main objects of interest for us because the concepts to be discussed in this paper are all about the local structure of derived schemes. Thus, it is enough to consider the affine case. More details will be given in §$\S$ 2.1.
Regarding certain geometric structures on higher spaces; such as $k$-shifted (closed) $p$-forms in the sense of [6], it is also known that for sufficiently “nice" cdgas (to be clear later), we can use the $A$-module $\Omega_{A}^{1}$ of Kähler differentials as a model for the cotangent complex $\mathbb{L}_{A}$ of $A$ so that we write $\mathbb{L}_{A}\simeq\Omega_{A}^{1}$. Then, by a $k$-shifted $p$-form on ${\rm Spec}A$ for $A$ a sufficiently nice cdga, we actually mean a $k$-cohomology class of the complex $(\Lambda^{p}\Omega_{A}^{1},d)$. Likewise, a $k$-shifted closed $p$-form on ${\rm Spec}A$ is just a $k$-cohomology class of the complex $\prod_{i\geq 0}\big{(}\Lambda^{p+i}\Omega^{1}_{A}[-i],\ d_{tot}=d+d_{dR}\big{)}$.
A reasonable notion of non-degeneracy is also available in this framework, which leads to the definition of a shifted symplectic structure. Loosely speaking, we are then able to define the notion of a $k$-shifted contact form on ${\rm Spec}A$ to be a $k$-shifted 1-form $\alpha$ on ${\rm Spec}A$ with the property that the $k$-shifted 2-form $d_{dR}\alpha$ satisfies a non-degeneracy condition, which will be formulated later. In fact, we will provide the general definition of contact data for derived (Artin) stacks; rather than just for affine derived ${\mathbb{K}}$-schemes with “nice" local models.
For shifted symplectic structures on derived schemes, it has been shown in [7] that for $k<0$, every $k$-shifted symplectic derived ${\mathbb{K}}$-scheme $(\bf X,\omega^{\prime})$ is Zariski locally equivalent to $({\rm Spec}A,\omega)$ with a pair $A,\omega$ in certain symplectic Darboux form. More precisely, for $k<0$, Bussi, Brav and Joyce [7, Theorem 5.18] proved that given a $k$-shifted symplectic derived ${\mathbb{K}}$-scheme $(\bf X,\omega^{\prime})$, one can find the so-called “minimal standard form" cdga $A$, a Zariski open inclusion $\iota:{\rm Spec}A\hookrightarrow\bf X$, and ``````` `coordinates" $x^{-i}_{j},y^{k+i}_{j}\in A$ with $\iota^{*}(\omega^{\prime})\sim(\omega^{0},0,0,\dots)$ such that
$\omega^{0}=\sum_{i,j}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}.$
We should point out that the expression of $\omega^{0}$ holds true only for the case where $k<0$ is an odd integer. The other possible cases require some modifications depending on whether $k/2$ is even or odd. However, the underlying idea behind the proofs for each case is the same.
Note also that the case $k<0$ odd is relatively simple and instructive enough to capture the essential techniques for the constructions of local models under consideration. Therefore, in this paper, we will mainly concentrate on the case with $k<0$ odd and use it as a prototype construction. For the other cases, we will not give all the details. Instead, we will only provide a brief outline. For details, we will always refer to [7, Examples 5.8, 5.9 & 5.10].
Results and the outline.
In this paper, we introduce the notion of a $k$-shifted contact structure on a derived (Artin) stack. The goals are then to develop Darboux-type models for negatively shifted contact structures and investigate further possible outcomes, such as symplectifications.
A brief definition. A $k$-shifted contact data will consist of a morphism $f:\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}$ of perfect complexes, a line bundle $L$ such that $Cone(f)\simeq L[k]$, and a locally defined $k$-shifted 1-form $\alpha:\mathbb{T}_{{\bf X}}\rightarrow\mathcal{O}_{{\bf X}}[k]$ satisfying a non-degeneracy condition (cf. Definitions 3.5 & 3.6). The next two theorems summarize the main results of this paper.
A Darboux-type theorem. We first discuss the existence of Darboux-type local models for $k$-shifted contact derived ${\mathbb{K}}$-schemes when $k<0$ (see Example 3.10). Next, for a locally finitely presented derived ${\mathbb{K}}$-scheme $\bf X$ with a $k$-shifted contact structure for $k<0,$ we prove the following result (cf. Theorem 3.13):
Theorem 1.1.
Every $k$-shifted contact derived ${\mathbb{K}}$-scheme $\bf X$ is locally equivalent to $({\rm Spec}A,\alpha_{0})$ for $A$ a minimal standard form cdga and $\alpha_{0}$ a contact Darboux form.
The symplectification. Secondly, we establish a shifted version of the classical connection between contact and symplectic geometries. In classical contact geometry, for a contact manifold $M$, there is a unique symplectified space with a symplectic structure canonically determined by the contact data of $M$. In this paper, we provide a similar result for shifted contact derived schemes.
The upshot is that given a (locally finitely presented) derived ${\mathbb{K}}$-scheme $\bf X$ with a $k$-shifted contact structure for $k<0,$ we define its symplectification $\mathcal{S}_{\bf{X}}$ of $\bf X$ as the total space of a certain $\mathbb{G}_{m}$-bundle over $\bf X$, constructed via the data of $k$-shifted contact structure, and provided with a canonical $k$-shifted symplectic structure on $\mathcal{S}_{\bf{X}}$ for which the $\mathbb{G}_{m}$-action is of weight 1 (cf. Definition 4.3 & Theorem 4.7). In brief, we have:
Theorem 1.2.
The space $\mathcal{S}_{\bf{X}}$ has the structure of a $k$-shifted symplectic derived stack with a symplectic form $\omega$ which is canonically determined by the shifted contact structure of $\bf X$. We then call the pair $(\mathcal{S}_{\bf{X}},\omega)$ the symplectification of $\bf X$.
Now, let us describe the content of this paper in more detail and provide an outline. In Section 2, we review derived symplectic geometry and symplectic Darboux forms. We begin by some background material on derived algebraic geometry and present nice local models for derived ${\mathbb{K}}$-schemes and their cotangent complexes. In §$\S$ 2.2, using these nice local models, we study shifted symplectic structures. §$\S$ 2.3 outlines symplectic Darboux forms on derived schemes and presents Darboux-type results given by Bussi, Brav and Joyce [7, Theorem 5.18].
Section 3 discusses contact structures. In §$\S$ 3.1, classical contact geometry is briefly revisited, and then, in §$\S$ 3.2, we introduce shifted contact structures and discuss their properties. In §$\S$ 3.3, we state a Darboux-type theorem for negatively shifted contact structures on derived ${\mathbb{K}}$-schemes (Theorem 3.13) and provide the proof of Theorem 1.1.
In Section 4, we discuss the concept of symplectification for (negatively) shifted contact derived ${\mathbb{K}}$-schemes and give the proof of Theorem 1.2 (cf. Definition 4.3 & Theorem 4.7).
Section 5 provides some concluding remarks on possible “stacky" generalizations of the main results of this paper. It also advertises possible future directions and some ongoing projects.
Acknowledgments.
I would like to thank Alberto Cattaneo and Ödül Tetik for useful discussions. It is also a pleasure to thank the Institute of Mathematics, University of Zurich, where this research was conducted. I personally benefited a lot from hospitality and research environment of the Institute.
I thank the anonymous Referees for the comprehensive review. I would like to express my gratitude to the Referees for the valuable comments and suggestions, which improved the quality of the manuscript.
The author acknowledges support of the Scientific and Technological Research Council of Turkey (TÜBİTAK) under 2219-International Postdoctoral Research Fellowship Program (2021-1).
Conventions.
Throughout the paper, $\mathbb{K}$ will be an algebraically closed field of characteristic zero. All cdgas will be graded in nonpositive degrees and over $\mathbb{K}.$ All classical ${\mathbb{K}}$-schemes will be locally of finite type, and all derived ${\mathbb{K}}$-schemes/Artin stacks X are assumed to be locally finitely presented.
2 Shifted symplectic structures
2.1 Some derived algebraic geometry
In this section, we outline the basics of DAG, present some material relevant to this paper, and state some useful results from shifted symplectic geometry. As stressed before, we use both Toën & Vezzosi’s version of DAG [4, 5] and the Lurie’s version [11]. In what follows, we just intend to give a brief sketch for the objects and constructions that we will be mostly interested in.
Definition 2.1.
Denote by $cdga_{{\mathbb{K}}}$ the category of commutative differential graded $\mathbb{K}$-algebras in non-positive degrees, where an object $A$ in $cdga_{{\mathbb{K}}}$ consists of
  1. (1)
    a collection of ${\mathbb{K}}$-vector spaces $\{A^{i}\}$, where $A^{i}$ is the ${\mathbb{K}}$-vector space of degree $i$ elements for $i=0,-1,\dots$,
  2. (2)
    a ${\mathbb{K}}$-bilinear, associative, supercommutative multiplication $A^{n}\otimes A^{m}\xrightarrow{\cdot}A^{n+m}$, and
  3. (3)
    a unique square-zero derivation of degree 1 (the differential) $\rm d$ on $A$ satisfying the graded Leibniz rule
    $\rm d(a\cdot b)=(\rm da)\cdot b+(-1)^{n}a\cdot(\rm db)$
    for all $a\in A^{n},b\in A^{m}.$
We denote such objects by $(A,\rm d)$ or just $A$. Note that $A$ has a decomposition $A=\bigoplus_{i}A^{i}.$
A morphism in $cdga_{{\mathbb{K}}}$, on the other hand, is a collection of degree-wise ${\mathbb{K}}$-linear morphisms $f=\{f^{i}\}:A\rightarrow B$ such that each $f^{i}:A^{i}\rightarrow B^{i}$ commutes with all the structures of $A,B$.
Definition 2.2.
Denote by $dSt_{\mathbb{K}}$ the $\infty$-category of derived stacks, where an object X of $dSt_{\mathbb{K}}$ is given as a certain $\infty$-functor
$\textbf{X}:\ cdga_{\mathbb{K}}\longrightarrow\ sSets,$ (2.1)
where $sSets$ denote the $\infty$-category of simplical sets. More precisely, objects in $dSt_{\mathbb{K}}$ are simplicial presheaves preserving weak equivalences and possessing the descent/local-to-global property w.r.t. the site structure on the source. For a brief review, we refer to [14].
We write $cdga_{{\mathbb{K}}}^{\infty}$ for the associated $\infty$-category of $cdga_{{\mathbb{K}}}$ such that the homotopy category $Ho(cdga_{{\mathbb{K}}}^{\infty})$ can be obtained from $cdga_{{\mathbb{K}}}$ by formally inverting quasi-isomorphsims.
Note that $Ho(cdga_{{\mathbb{K}}}^{\infty})$ is just an ordinary category. We should also point out that $cdga_{{\mathbb{K}}}^{\infty}$, $cdga_{{\mathbb{K}}}$ and $Ho(cdga_{{\mathbb{K}}}^{\infty})$ have the same objects; however, lifting properties of morphisms are different. That is, a morphism $f:A\rightarrow B$ in $cdga_{{\mathbb{K}}}$ is also a morphism in $cdga_{{\mathbb{K}}}^{\infty}$ and $Ho(cdga_{{\mathbb{K}}}^{\infty})$. But in general, the converse is not true unless $A$ is cofibrant. In the rest of this paper, we will be interested in certain types of cdgas, called standard form cdgas, which are in fact “sufficiently cofibrant", and hence suitable for our purposes.
In this framework, there also exists an appropriate concept of a spectrum functor [11, §$\S$ 4.3]
${\rm Spec}:(cdga_{{\mathbb{K}}}^{\infty})^{op}\rightarrow dSt_{{\mathbb{K}}},$
which leads to the following definition.
Definition 2.3.
An object X in $dSt_{\mathbb{K}}$ is called an affine derived $\mathbb{K}$-scheme if $\textbf{X}\simeq{\rm Spec}A$ for some cdga $A\in cdga_{{\mathbb{K}}}$. An object $\bf X$ in $dSt_{\mathbb{K}}$ is then called a derived $\mathbb{K}$-scheme if it can be covered by Zariski open affine derived $\mathbb{K}$-schemes $Y\subset X$.
Denote by $dSch_{{\mathbb{K}}}\subset dSt_{\mathbb{K}}$ the full $\infty$-subcategory of derived $\mathbb{K}$-schemes, and we simply write $dAff_{{\mathbb{K}}}\subset dSch_{\mathbb{K}}$ for the full $\infty$-subcategory of affine derived $\mathbb{K}$-schemes. Note that ${\rm Spec}:(cdga_{{\mathbb{K}}}^{\infty})^{op}\rightarrow dAff_{{\mathbb{K}}}$ gives an equivalence of $\infty$-categories.
We should note that throughout this paper, $\mathbb{K}$ will be an algebraically closed field of characteristic zero. We also assume that all classical ${\mathbb{K}}$-schemes are locally of finite type, and all derived ${\mathbb{K}}$-schemes X are locally finitely presented, by which we mean that X can be covered by Zariski open affines ${\rm Spec}A,$ where $A$ is a finetely presented cdga over ${\mathbb{K}}.$
Remark 2.4.
Thanks to the Yoneda embedding, one can also realize algebro-geometric objects (like classical ${\mathbb{K}}$-schemes, stacks, derived "spaces", etc…) as certain functors in addition to the standard ringed-space formulation. We have the following enlightening diagram from [14] encoding such a functorial interpretation:
[…]
Here $CAlg_{\mathbb{K}}$ denotes the category of commutative ${\mathbb{K}}$-algebras. Denote by $St_{{\mathbb{K}}}$ the $\infty$-category of (higher) ${\mathbb{K}}$-stacks, where objects in $St_{{\mathbb{K}}}$ are defined via the diagram above.
In the underived setup, we have the classical “spectrum functor"
${\rm spec}:(CAlg_{\mathbb{K}})^{op}\rightarrow St_{{\mathbb{K}}}.$
We then call an object $X$ of $St_{{\mathbb{K}}}$ an affine ${\mathbb{K}}$-scheme if $X\simeq{\rm spec}A$ for some $A\in CAlg_{\mathbb{K}}$, and a ${\mathbb{K}}$-scheme if it has an open cover by affine ${\mathbb{K}}$-schemes.
In addition to the spectrum functors ${\rm Spec},{\rm spec}$ above, there is a natural truncation functor $\tau:dSt_{{\mathbb{K}}}\rightarrow St_{{\mathbb{K}}}$, along with a fully faithfull left adjoint inclusion functor $\iota:St_{{\mathbb{K}}}\hookrightarrow dSt_{{\mathbb{K}}}$, which can be thought of as an embedding of classical algebraic ${\mathbb{K}}$-spaces into derived spaces.
Note that, for a cdga $A$ there exists an equivalence $\tau\circ{\rm Spec}A\simeq{\rm spec}H^{0}(A)$. This means that if $\bf X$ is a (affine) derived ${\mathbb{K}}$-scheme, then its truncation $=\tau(\textbf{X})$ X ) is a (affine) ${\mathbb{K}}$-scheme. Therefore, we can consider a derived ${\mathbb{K}}$-scheme $\bf X$ as an infinitesimal thickening of its truncation $X$. It follows that points of a derived ${\mathbb{K}}$-scheme $\bf X$ are the same as points of of its truncation $X$. It means that the main difference between $X$ and $\bf X$ is in fact encoded by the scheme structure, not by the points!
Nice local models for derived ${\mathbb{K}}$-schemes.
The following result (Theorem 2.5) plays an important role in constructing useful local algebraic models for derived ${\mathbb{K}}$-schemes and for the so-called $k$-shifted symplectic structures on them.
The upshot is that given a derived ${\mathbb{K}}$-scheme $\bf X$ (locally of finite presentation) and a point $x\in\bf X$, one can always find a “refined" local affine neighborhood ${\rm Spec}A$ of $x$ that allows us to make more explicit computations over this neighborhood. For example, using such local models, we can identify the cotangent complex $\mathbb{L}_{A}$ with the module of Kähler differentials $\Omega_{A}^{1}$, and then we can provide explicit representatives (rather than just cohomology classes) for (closed) $p$-forms of degree $k$. In this regard, Bussi, Brav and Joyce proved the following theorem.
Theorem 2.5.
([7, Theorem 4.1]) Every derived $\mathbb{K}$-scheme $X$ is Zariski locally modelled on ${\rm Spec}A$ for some “minimal standard form" cdga $A$ in $cdga_{\mathbb{K}}$.
More precisely, for each $x\in X$ there is a pair $\big{(}A,i:{\rm Spec}A\hookrightarrow X\big{)}$ and $p\in{\rm Spec}H^{0}(A)$ such that $i$ is an open inclusion with $i(p)=x$, where $A$ is a special kind of cdga (cf. Definition 2.6).
Moreover, there is a reasonable way to compare two such local charts $i:{\rm Spec}A\hookrightarrow X$ and $j:{\rm Spec}B\hookrightarrow X$ on their overlaps via a third chart. For details, see [7, Theorem 4.1 & 4.2].
In the remainder of this section, we shall elaborate the content of Theorem 2.5, and introduce appropriate notions for the constructions of interest. We will closely follow [7, 8].
Definition 2.6.
$A\in cdga_{\mathbb{K}}$ is of standard form if $A^{0}$ is a smooth finitely generated $\mathbb{K}$-algebra, the module $\Omega^{1}_{A^{0}}$ of Kähler differentials is free $A^{0}$-module of finite rank, and the graded algebra $A$ is freely generated over $A^{0}$ by finitely many generators, all in negative degrees.
In fact, there is a systematic way of constructing such cdgas. [7, Example 2.8] explains how to build these cdgas starting from a smooth ${\mathbb{K}}$-algebra $A^{0}:=A(0)$ via applying a sequence of localizations. The upshot is follows: Let $n\in{\mathbb{N}}$, then a cdga $A$, as a commutative graded algebra, can be constructed inductively from a smooth ${\mathbb{K}}$-algebra $A(0)$ by adjoining free finite rank modules $M^{-i}$ of generators in degree $-i$ for $i=1,2,\dots,n.$ More precisely, for any given $n\in{\mathbb{N}}$, we can inductively construct a sequence of cdgas
$A(0)\rightarrow A(1)\rightarrow\cdots\rightarrow A(i)\rightarrow\cdots A(n)=:A,$ (2.2)
where $A^{0}:=A(0)$, and $A(i)$ is obtained from $A(i-1)$ by adjoining generators in degree $-i$, given by $M^{-i}$, for all $i$. Here, each $M^{-i}$ is a free finite rank module (of degree $-i$ generators) over $A(i-1)$. Therefore, the underlying commutative graded algebra of $A=A(n)$ is freely generated over $A(0)$ by finitely many generators, all in negative degrees $-1,-2,\dots,-n$.
Definition 2.7.
A standard form cdga $A$ is said to be minimal at $p\in{\rm Spec}H^{0}(A)$ if $A=A(n)$ is defined by using the minimal possible numbers of graded generators in each degree $\leq 0$ compared to all other cdgas locally equivalent to $A$ near $p$. (There will be an equivalent definition for minimality later, see Definition 2.16.)
Definition 2.8.
Let $A$ be a standard form cdga. $A^{\prime}\in cdga_{\mathbb{K}}$ is called a localization of $A$ if $A^{\prime}$ is obtained from $A$ by inverting an element $f\in A^{0}$, by which we mean $A^{\prime}=A\otimes_{A^{0}}A^{0}[f^{-1}]$.
$A^{\prime}$ is then of standard form with $A^{\prime 0}\simeq A^{0}[f]$. If $p\in{\rm spec}H^{0}(A)$ with $f(p)\neq 0$, we say $A^{\prime}$ is a localization of $A$ at $p$.
With these definitions in hand, one has the following observations:
Observation 2.9.
Let $A$ be a standard form cdga. If $A^{\prime}$ is a localization of $A$, then ${\rm Spec}A^{\prime}\subset{\rm Spec}A$ is a Zariski open subset. Likewise, if $A^{\prime}$ is a localization of $A$ at $p\in{\rm spec}H^{0}(A)\simeq\tau({\rm Spec}A)$, then ${\rm Spec}A^{\prime}\subset{\rm Spec}A$ is a Zariski open neighborhood of $p.$
Observation 2.10.
Let $A=A(k)$ be a standard form cdga, then there exist generators $x_{1}^{-i},x_{2}^{-i},\cdots,x_{m_{i}}^{-i}$ in $A^{-i}$ (after localization, if necessary) with $i=1,2,\cdots,k$  and $m_{i}\in\mathbb{Z}_{\geq 0}$ such that
$A=A(0)\big{[}x_{j}^{-i}:i=1,2,\dots,k,\ j=1,2,\dots,m_{i}\big{]},$ (2.3)
where the subscript $j$ in $x_{j}^{i}$ labels the generators, and the superscript $i$ indicates the degree of the corresponding element. So, we can consider $A$ as a graded polynomial algebra over $A(0)$ on finitely many generators, all in negative degrees.
Definition 2.11.
We then define the virtual dimension of $A$ to be the integer ${\rm vdim\ }A=\sum_{i}(-1)^{i}m_{i}$.
Observation 2.12.
Geometrically, the “smoothness"""" condition on $A^{0}$ implies that the corresponding affine $\mathbb{K}$-scheme $U={\rm spec}A^{0}$ is smooth together with a local ($\acute{e}tale$) coordinate system
$(x_{1}^{0},x_{2}^{0},\dots,x_{m_{0}}^{0}):U\longrightarrow\mathbb{A}^{m_{0}}_{\mathbb{K}}.$ (2.4)
Nice local models for cotangent complexes of derived schemes.
Given $A\in cdga_{\mathbb{K}}$, $d$ on $A$ induces a differential on $\Omega_{A}^{1}$, denoted again by $d$. This makes $\Omega_{A}^{1}$ into a $\rm dg$-module $(\Omega_{A}^{1},d)$ with the property that $\delta\circ d=d\circ\delta,$ where $\delta:A\rightarrow\Omega_{A}^{1}$ is the universal derivation of degree 00.
Write the decomposition of $\Omega^{1}_{A}$ into graded pieces $\Omega^{1}_{A}=\bigoplus_{k=-\infty}^{0}\big{(}\Omega^{1}_{A}\big{)}^{k}$ with the differential $d:\big{(}\Omega^{1}_{A}\big{)}^{k}\longrightarrow\big{(}\Omega^{1}_{A}\big{)}^{k+1}$. Then we define the de Rham algebra of $A$ as a double complex
$DR(A)=Sym_{A}(\Omega_{A}^{1}[1])\simeq\bigoplus\limits_{p=0}^{\infty}\bigoplus\limits_{k=-\infty}^{0}\big{(}\Lambda^{p}\Omega^{1}_{A}\big{)}^{k}[p],$ (2.5)
where $DR(A)$ has two gradings: the grading w.r.t. $p$ is called the weight, and the grading w.r.t. $k$ is called the degree. By construction, there are two differentials, namely the internal differential d and the de Rham differential $d_{dR}.$ We diagrammatically have
[…]
such that $d_{tot}=d+d_{dR}$, and both differentials satisfy the relations
$d^{2}=d_{dR}^{2}=0,\text{\ and \ }d\circ d_{dR}+d_{dR}\circ d=0.$ (2.7)
We also have the natural multiplication on $DR(A)$:
$\big{(}\Lambda^{p}\Omega^{1}_{A}\big{)}^{k}[p]\times\big{(}\Lambda^{q}\Omega^{1}_{A}\big{)}^{\ell}[q]\longrightarrow\big{(}\Lambda^{p+q}\Omega^{1}_{A}\big{)}^{k+\ell}[p+q].$ (2.8)
Observation 2.13.
It should be noted that the constructions of $\Omega_{A}^{1}\text{ and }DR(A)$ depend only on the underlying commutative graded algebra of $A$, not on the differential $d$ on $A.$
Remark 2.14.
When $A=A(k)$ is a minimal standard form cdga, there are two important outcomes:
  1. 1.
    With such local coordinates $(x_{1}^{0},x_{2}^{0},\cdots,x_{m_{0}}^{0})$, we have
    $\Omega^{1}_{A^{0}}\cong A^{0}\otimes_{\mathbb{K}}\langle d_{dR}x_{1}^{0},\cdots,d_{dR}x_{m_{0}}^{0}\rangle_{\mathbb{K}}.$ (2.9)
    Furthermore, the Kähler differentials is a $A$-module of the form
    $\Omega^{1}_{A}\cong A\otimes_{\mathbb{K}}\langle d_{dR}x_{j}^{-i}:i=0,1,2,\cdots,k,\ j=1,2,\cdots,m_{i}\rangle_{\mathbb{K}}.$ (2.10)
  2. 2.
    $\Omega^{1}_{A}$ provides a local model for the cotangent complex $\mathbb{L}_{A}$. That is, in the case of a minimal standard form cdga, the cotangent complex $\mathbb{L}_{A}$ has the identification
    $\mathbb{L}_{A}=\Omega^{1}_{A}.$ (2.11)
Note that if $D(Mod_{A})$ denotes the derived category of $Mod_{A}$, then $\mathbb{L}_{A}\in D(Mod_{A})$ for standard form cdgas. In general, even if both $\mathbb{L}_{A}\text{ and }\Omega^{1}_{A}$ are closely related, the identification in (2.11) is not true for an arbitrary $A\in cdga_{\mathbb{K}}$ [7].
When $A=A(n)$ is a standard form cdga as in (2.2), we also have the following description for the restriction of the cotangent complex $\mathbb{L}_{A}$ to ${\rm spec}H^{0}(A)$.
Proposition 2.15.
([7, Prop. 2.12]) If $A=A(n)$, with $n\in{\mathbb{N}}$, is a standard form cdga constructed inductively as in (2.2), then the restriction of $\mathbb{L}_{A}$ to ${\rm spec}H^{0}(A)$ is represented by a complex of $H^{0}(A)$-modules
$0\rightarrow V^{-n}\xrightarrow{d^{-n}}V^{-n+1}\rightarrow\cdots\rightarrow V^{-1}\xrightarrow{d^{-1}}V^{0}\rightarrow 0,$ (2.12)
where each $V^{-i}$ can in fact be defined as $V^{-i}=H^{-i}(\mathbb{L}_{A(i)/A(i-1)})$, with $\mathbb{L}_{A(i)/A(i-1)}$ the relative cotangent complex of the map $A(i-1)\rightarrow A(i)$ in (2.2) satisfying
$\mathbb{L}_{A(i)/A(i-1)}\simeq A(i)\otimes_{A(i-1)}M^{-i}[i].$
Moreover, the differential $V^{-i}\xrightarrow{d^{-i}}V^{-i+1}$ is identified with the composition
$H^{-i}(\mathbb{L}_{A(i)/A(i-1)})\rightarrow H^{-i+1}(\mathbb{L}_{A(i-1)})\rightarrow H^{-i+1}(\mathbb{L}_{A(i-1)/A(i-2)}),$
which can be obtained from the fiber sequences induced by the maps $A(i-1)\rightarrow A(i)$ in (2.2). Note in particular that $j>-i,$ we have $H^{j}(\mathbb{L}_{A(i)/A(i-1)})=0.$ More details and the proof can be found in [7, Prop. 2.12].
With this result in hand, using local coordinates above, write
$V^{-i}=\langle d_{dR}x^{-i}_{1},d_{dR}x^{-i}_{2},\dots,d_{dR}x^{-i}_{m_{i}}\rangle_{A(0)}\ \text{ for }i=0,1,\dots,n.$
It follows that we have a similar local description for the tangent complex $\mathbb{T}_{A}=(\mathbb{L}_{A})^{\vee}$ of $A$ when restricted to ${\rm spec}H^{0}(A)$. Also, we have an alternative definition of minimality at a point $p\in{\rm spec}H^{0}(A)$ for a cdga of the form $A=A(n)$.
Definition 2.16.
Let $A=A(n)$, with $n\in{\mathbb{N}}$, be a standard form cdga constructed inductively as in (2.2). $A$ is said to be minimal at $p\in{\rm spec}H^{0}(A)$ if the internal differential $d^{-i}|_{p}=0$ in the complex $\mathbb{L}_{A}|_{{\rm spec}H^{0}(A)}$ given in (2.12).
Note from Proposition 2.15 that at $p\in{\rm spec}H^{0}(A)$, with $A$ a standard form cdga, $\Omega^{1}_{A}|_{p}$ is a complex of ${\mathbb{K}}$-vector spaces and that Definition 2.16 implies $m_{i}=\dim(H^{-i}(\mathbb{L}_{A}|_{p}))$ for each $i$, and hence $A$ is defined by using the minimum number of graded variables in each degree $\leq 0$ compared to all other cdgas locally equivalent to $A$ near $p$. Therefore, one recovers Definition 2.7.
2.2 PTVV’s shifted symplectic geometry on derived schemes
Let $\bf X$ be a locally finitely presented derived ${\mathbb{K}}$-scheme with $p\geq 0,\ k\in{\mathbb{Z}}$. Pantev et al. [6] define simplicial sets of $p$-forms of degree $k$ and closed $p$-forms of degree $k$ on $\bf X.$ Denote these simplicial sets by $\mathcal{A}^{p}(X,k)$ and $\mathcal{A}^{(p,cl)}(X,k)$, respectively. These definitions are in fact given first for affine derived ${\mathbb{K}}$-schemes. Later, both concepts are defined for a general $\bf X$ in terms of mapping stacks. A summary of key ideas can be found in [7, §$\S$ 3.4]
In our case, we consider $\textbf{X}={\rm Spec}A$ with $A$ a standard form cdga, and hence take $\Lambda^{p}\mathbb{L}_{A}=\Lambda^{p}\Omega^{1}_{A}$. Therefore, elements of $\mathcal{A}^{p}(X,k)$ form a simplicial set such that $k$-cohomology classes of the complex $\big{(}\Lambda^{p}\Omega^{1}_{A},d\big{)}$ correspond to the connected components of this simplicial set. Likewise, the connected components of $\mathcal{A}^{(p,cl)}(X,k)$ are identified with the $k$-cohomology classes of the complex $\prod_{i\geq 0}\big{(}\Lambda^{p+i}\Omega^{1}_{A}[-i],d_{tot}\big{)}$. We want to work with explicit representatives for these classes.
It should be noted that the results that are cited or to be proven in this paper are all about the local structure of derived schemes. Thus, it is enough to consider the affine case. Moreover, we always assume all local models are sufficiently nice by using Theorem 2.5 if necessary.
Definition 2.17.
Let $\textbf{X}={\rm Spec}A$ be an affine derived $\mathbb{K}$-scheme for $A$ a minimal standard form cdga. A $p$-form of degree $k$ on $\bf X$ for $p\geq 0$ and $k\leq 0$ is an element
$\omega^{0}\in\big{(}\Lambda^{p}\Omega^{1}_{A}\big{)}^{k}\text{ with }d\omega^{0}=0.$ (2.13)
Note that an element $\omega^{0}$ defines a cohomology class as being $d$-closed. That is,
$[\omega^{0}]\in H^{k}\big{(}\Lambda^{p}\Omega^{1}_{A},d\big{)},$
where two $p$-forms $\omega_{1}^{0},\omega_{2}^{0}$ of degrees $k$ are equivalent if there exists $\alpha_{1,2}\in\big{(}\Lambda^{p}\Omega^{1}_{A}\big{)}^{k-1}$ so that $\omega_{1}^{0}-\omega_{2}^{0}=d\alpha_{1,2}$.
Remark 2.18.
In the classical ``````` `underived"""" case, for instance when $X={\rm spec}A$ is smooth for a commutative $\mathbb{K}$-algebra $A$, the cotangent complex $\mathbb{L}_{X}$ is just a vector bundle over $X$, and denoted simply by $T^{*}X$. Then, a $p$-form $\omega$ on $X$ is defined to be a global section of the bundle $\Lambda^{p}T^{*}X$. A careful observation reveals that Definition 2.17 does generalize the definition of a $p$-form on a smooth space in the following sense: It is clear that any commutative $\mathbb{K}$-algebra $A$ can be realized as an object in $cdga_{\mathbb{K}}$ concentrated in degree 0 with the trivial differential. Thus, in the language of Definition 2.17, a naïve notion of ``````` `a $p$-form $\omega$ on a smooth space $X"$ is just a $p$-form $\omega$ of degree 0 on $X\simeq\tau\circ\iota(X)$ in $St_{{\mathbb{K}}}$ such that $\omega\in\big{(}\Lambda^{p}T^{*}X\big{)}^{0}.$ Note that the condition $d\omega=0$ holds trivially, and hence $[\omega]\in H^{0}\big{(}\Lambda^{p}T^{*}X,d=0\big{)}$. Here, $\mathbb{L}_{X}=T^{*}X$ is again viewed as graded object concentrated in degree 0, with the zero differential.
In DAG, on the other hand, $\Lambda^{p}\mathbb{L}_{X}$ is a (double) complex which possesses a non-trivial internal differential as above, and hence one needs to take into account higher non-trivial cohomology groups as well.
Definition 2.19.
A 2-form $\omega^{0}$ of degree $k$ on ${\rm Spec}A$ for $A$ a minimal standard form cdga is called non-degenerate if the induced morphism $\omega^{0}:\mathbb{T}_{A}\rightarrow\Omega^{1}_{A}[k],\ Y\mapsto\iota_{Y}\omega^{0}$, is a quasi-isomorphism, where $\mathbb{T}_{A}=(\mathbb{L}_{A})^{\vee}={\rm Hom}_{A}(\Omega^{1}_{A},A)$ is the tangent complex of $A$.
Definition 2.20.
Let $\textbf{X}={\rm Spec}A$ be an affine derived $\mathbb{K}$-scheme with $A$ a minimal standard form cdga. A closed $p$-form of degree $k$ on $\bf X$ for $p\geq 0$ and $k\leq 0$ is a sequence $\omega=(\omega^{0},\omega^{1},\dots)$ with $\omega^{i}\in\big{(}\Lambda^{p+i}\Omega^{1}_{A}\big{)}^{k-i}$ satisfying the following conditions:
  1. (1)
    $d\omega^{0}=0$ in $\big{(}\Lambda^{p}\Omega^{1}_{A}\big{)}^{k+1}.$
  2. (2)
    $d_{dR}\omega^{i}+d\omega^{i+1}=0$ in $\big{(}\Lambda^{p+i+1}\Omega^{1}_{A}\big{)}^{k-i}$, $i\geq 0.$
Remark 2.21.
  1. 1.
    From Definition 2.20, there exists a natural projection morphism
    $\pi:\mathcal{A}^{(p,cl)}(X,k)\longrightarrow\mathcal{A}^{p}(X,k),\ \ \omega=(\omega^{i})_{i\geq 0}\longmapsto\omega^{0}.$ (2.14)
  2. 2.
    When we restrict ourselves to the classical case as in Remark 2.18, the one in which everything is concentrated in degree 0, we have $d=0$ and hence $d_{tot}=d_{dR}$. Moreover, the only possible non-trivial component of $\omega$ is $\omega^{0}$. Therefore, using the truncation functor as before, the conditions in Definition 2.20 reduce to
    $\omega^{0}\in H^{0}\big{(}\Lambda^{p}T^{*}X\big{)}\text{ with }d_{dR}\omega^{0}=0.$ (2.15)
    Thus, Definition 2.20 reduces to the usual definition of a (de Rham) closed $p$-form on smooth spaces.
Definition 2.22.
A closed $2$-form $\omega=(\omega^{i})_{i\geq 0}$ of degree $k$ on an affine derived $\mathbb{K}$-scheme ${\rm Spec}A$ for a minimal standard form cdga $A$ is called a $k$-shifted symplectic structure if $\pi(\omega)=\omega^{0}$ is a non-degenerate 2-form of degree $k$.
2.3 Shifted symplectic Darboux models
One of the main theorems in [7] provides a $k$-shifted version of the classical Darboux theorem in symplectic geometry. The statement is as follows.
Theorem 2.23.
([7, Theorem 5.18]) Given a derived $\mathbb{K}$-scheme $X$ with a $k$-shifted symplectic form $\omega^{\prime}$ for $k<0$ and $x\in X$, there is a local model $\big{(}A,f:specA\hookrightarrow X,\omega\big{)}$ and $p\in{\rm Spec}H^{0}(A)$ such that $f$ is an open inclusion with $f(p)=x$, $A$ is a standard form that is minimal at $p$, and $\omega$ is a $k$-shifted symplectic form on ${\rm Spec}A$ such that $A,\omega$ are in Darboux form, and $f^{*}(\omega^{\prime})\sim\omega$ in the space of $k$-shifted closed 2-forms.
To be more precise, it is proven in [7, Theorem 5.18] that such $\omega$ can be constructed explicitly depending on the integer $k<0$. Indeed, there are three cases in total:
$(1)\ k\text{ is odd},\ (2)\ k\equiv 0\mod 4,\ (3)\ k\equiv 2\mod 4.$
Equivalently, the cases can be expressed as $(1)\ k/2\notin{\mathbb{Z}},\ (2)\ k/2\text{ is even},\text{ and }(3)\ k/2\text{ is odd},$ respectively. In short, Theorem 2.23 says that every $k$-shifted symlectic derived ${\mathbb{K}}$-scheme $(\bf X,\omega^{\prime})$ is Zariski locally equivalent to $({\rm Spec}A,\omega)$ for some $A,\omega$, where $A$ is a minimal standard form cdga and $\omega$ is a $k$-shifted symplectic form on ${\rm Spec}A$ such that $\omega$ is given in a standard way depending on the cases above.
In this paper, for simplicity, we will examine a family of explicit Darboux forms for $k<0$ an odd integer [7, Example 5.8]. The other cases can be studied in a similar way, but with some modifications. We will outline the steps. More details can be found in [7, §$\S$ 5.3].
We first begin with a useful result that plays a significant role in constructing Darboux-type local models below. The upshot is that one can always simplify the form of a given closed 2-form $\omega=(\omega^{0},\omega^{1},\omega^{2},\dots)$ of degree $k<0$ on ${\rm Spec}A$ so that $\omega^{0}$ can be taken to be exact and $\omega^{i}=0$ for all $i>0.$ More precisely, we have the following result.
Proposition 2.24.
([7, Prop. 5.7] ) Let $\omega=(\omega^{0},\omega^{1},\omega^{2},\dots)$ be a closed 2-form of degree $k<0$ on ${\rm Spec}A$ for $A$ a standard form cdga over ${\mathbb{K}}.$ Then there exist $H\in A^{k+1}$ and $\phi\in(\Omega^{1}_{A})^{k}$ such that $dH=0$ in $A^{k+2}$, $d_{dR}H+d\phi=0$ in $(\Omega^{1}_{A})^{k+1}$, and $\omega\sim(d_{dR}\phi,0,0,\dots).$
Moreover, if $(H^{\prime},\phi^{\prime})$ is another such pair for fixed $\omega,k,A,$ then there exist $h\in A^{k}$ and $\sigma\in(\Omega_{A}^{1})^{k-1}$ such that $H-H^{\prime}=dh$ and $\phi-\phi^{\prime}=d_{dR}h+d\sigma.$
The proof of Proposition 2.24 is based on the fact that any such forms can be interpreted in the context of cyclic homology theory of mixed complexes. Indeed, any such forms can be viewed as cocycles in the so-called negative cyclic complex of weight p on ${\rm Spec}A$, which is constructed from the de Rham algebra $DR(A)$ in certain way. When $p=2$, there are some useful short exact sequences and vanishing results, by which one can eventually obtain the desired simplification above. For more details on this cyclic homology perspective, we refer to [7, §$\S$ 5.2].
Observation 2.25.
Assume $(H,\phi)$ is a such pair for fixed $\omega,k,A,$ with $d_{dR}\phi=k\omega^{0}.$ Let $f\in{\mathbb{K}}$ be a non-zero element. Define $H^{\prime}=fH$ and $\phi^{\prime}=f\phi$. Then both $H^{\prime},\phi^{\prime}$ satisfy the relations $dH^{\prime}=0$ and $d_{dR}H^{\prime}+d\phi^{\prime}=0$. From the choices, we also have $d_{dR}\phi^{\prime}=kf\omega^{0},$ and hence $\omega\sim(d_{dR}\phi^{\prime},0,0,\dots).$ By Proposition 2.24, there exist $h\in A^{k}$ and $\sigma\in(\Omega_{A}^{1})^{k-1}$ such that $H-H^{\prime}=dh$ and $\phi-\phi^{\prime}=d_{dR}h+d\sigma.$ It follows that $(1-f)H=dh$. Localizing $A$ by the element $(1-f)$ if necessary, we can write $H=d\big{[}(1-f)^{-1}h\big{]}.$ It means that we can ``$locally"$ take $H$ to be $d$-exact.
Prototype Darboux model for $k<0$ odd.
Let $k=-2\ell-1$ for $\ell\in{\mathbb{N}}$. Then the local model consists of the following data:
  1. 1.
    Let $A^{0}=A(0)$ be a smooth $\mathbb{K}$-algebra of $\dim m_{0}$, choose $x_{1}^{0},\dots,x_{m_{0}}^{0}$ such that $d_{dR}x_{1}^{0},\dots,d_{dR}x_{m_{0}}^{0}$ form a basis for $\Omega_{A^{0}}^{1}$. Then $A$ is defined to be the free graded algebra over $A^{0}$ generated by variables
[…]
  2. 2.
    $\Omega^{1}_{A}$ is the free $A$-module of finite rank given by
    $\Omega^{1}_{A}\simeq A\otimes_{\mathbb{K}}\langle d_{dR}x_{j}^{-i},d_{dR}y_{j}^{k+i}:i=0,1,\dots,\ell,\ j=1,2,\cdots,m_{i}\rangle_{\mathbb{K}}.$ (2.17)
  3. 3.
    Define an element $\omega^{0}\in(\Lambda^{2}\Omega^{1}_{A})^{k}$ of degree $k$ and weight 2 in $DR(A)$ to be
    $\omega^{0}=\displaystyle\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}.$ (2.18)
  4. 4.
    It follows from Proposition 2.24 that there exists a pair $(\phi,H)\in(\Omega^{1}_{A})^{k}\times A^{k+1}$ satisfying the following properties:
    1. (a)
      $dH=0$ in $A^{k+2}$,  $d_{dR}H+d\phi=0$ in $(\Omega^{1}_{A})^{k+1}$, and $d_{dR}\phi=k\omega^{0}.$
    2. (b)
      $H$ satisfies the condition (a.k.a. the classical master equation)
      $\sum_{i=1}^{\ell}\sum_{j=1}^{m_{i}}\dfrac{\partial H}{\partial x_{j}^{-i}}\dfrac{\partial H}{\partial y_{j}^{k+i}}=0\text{ in }A^{k+2},$ (2.19)
      which in fact corresponds to the condition ``$dH=0".$ We call $H$ the Hamiltonian.
    3. (c)
      Explicitly,  we have
      $\phi:=\displaystyle\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}\big{[}-ix_{j}^{-i}d_{dR}y_{j}^{k+i}+(k+i)y_{j}^{k+i}d_{dR}x_{j}^{i}\big{]}.$ (2.20)
      Note that we can choose another representatives by replacing $H,\phi$ by $\phi^{\prime}=\phi+d_{dR}\theta$ and $H^{\prime}=H+d\theta$ for any $\theta\in A^{k}$. This modification will leave $\omega^{0}$ unchanged, and both $H^{\prime},\phi^{\prime}$ satisfy $dH^{\prime}=0$ and $d_{dR}H^{\prime}+d\phi^{\prime}=0$. Letting $\theta=\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}\big{[}(-1)^{i}x_{j}^{-i}y_{j}^{k+i}\big{]}$, for instance, we may take $\phi:=k\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}y_{j}^{k+i}d_{dR}x_{j}^{-i}$.
    4. (d)
      The internal differential $d$ on $A$ can be defined as
      $d|_{A^{0}}=0,\ \ dx_{j}^{-i}=\dfrac{\partial H}{\partial y_{j}^{k+i}}\ \text{ and }\ dy_{j}^{k+i}=\dfrac{\partial H}{\partial x_{j}^{-i}}.$ (2.21)
  5. 5.
    Clearly $d_{dR}\omega^{0}=0$, but it is a little bit cumbersome to check that $d\omega^{0}=0$, and $\omega^{0}$ defines a non-degenerate pairing. For details, we refer to [7, Example 5.8]. As a result, the sequence $\omega:=(\omega^{0},0,0,\dots)$ defines a $k$-shifted symplectic structure on ${\rm Spec}A$.
Definition 2.26.
If $A,\omega$ are as above, then we say that the pair $(A,\omega)$ is in (symplectic) Darboux form.
In brief, Theorem 2.23 implies that every $k$-shifted symlectic derived ${\mathbb{K}}$-scheme $(\bf X,\omega^{\prime})$, with $k<0$ odd, is Zariski locally equivalent to $({\rm Spec}A,\omega)$ for a pair $A,\omega$ in Darboux form as above. More precisely, Bussi, Brav and Joyce [7, Theorem 5.18] proved that given a $k$-shifted symlectic derived ${\mathbb{K}}$-scheme $(\bf X,\omega^{\prime})$, one can find a minimal standard form cdga $A$ with ``````` `coordinates" $x^{-i}_{j},y^{k+i}_{j}$, and a Zariski open inclusion $\iota:{\rm Spec}A\hookrightarrow\bf X$ such that $\iota^{*}(\omega^{\prime})\simeq(\omega^{0},0,0,\dots)$ and $\omega^{0}=\sum_{i,j}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}.$
Note that the expression of $\omega^{0}$ above is valid only for the case $k<0$ odd, and the other cases require some modifications and extra variables depending on whether $k/2$ is even or odd. But, as mentioned before, the proofs follow the same logic. We now give an outline for the cases.
Darboux forms for the other cases of $k.$
For the sake of completeness, we briefly summarize the cases when $k/2$ is even or odd. Here, the main difference from the case $k$ being odd is about the existence of middle degree variables. In fact, when $k$ is odd, there is no such degree. But if $k/2$ is even, there are such variables and 2-forms are anti-symmetric in these variables. On the other hand, when $k/2$ is odd, such forms are symmetric in the middle degree variables. Let us briefly examine each case:
  1. (a)
    [7, Example 5.9] When $k/2$ is even, say $k=-4\ell$ for $\ell\in{\mathbb{N}}$, the cdga $A$ is now free over $A(0)$ generated by the new set of variables
    $x_{1}^{-i},x_{2}^{-i},\dots,x_{m_{i}}^{-i}\text{ in degree }-i\text{ for }i=1,2,\dots,2\ell-1,$
    $x_{1}^{-2\ell},x_{2}^{-2\ell},\dots,x_{m_{2\ell}}^{-2\ell},y_{1}^{-2\ell},y_{2}^{-2\ell},\dots,y_{m_{2\ell}}^{-2\ell}\text{ in degree }-2\ell,$
    $y_{1}^{k+i},y_{2}^{k+i},\dots,y_{m_{i}}^{k+i}\text{ in degree }k+i\text{ for }i=0,1,\dots,2\ell-1.$ (2.22)
    Then we define an element $\omega^{0}=\sum_{i=0}^{2\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}$ in $(\Lambda^{2}\Omega^{1}_{A})^{k}$, and set $\omega$ to be $(\omega^{0},0,0,\dots)$ as before. Choose an element $H\in A^{k+1}$, the Hamiltonian, satisfying the analogue of classical master equation, and define $d$ on $A$ as in Equation (2.21) using $H$. We also define the element $\phi\in(\Omega^{1}_{A})^{k}$ by the analogue of Equation (2.20).
  2. (b)
    [7, Example 5.10] When $k/2$ is odd, say $k=-4\ell-2$ for $\ell\in{\mathbb{N}}$, $A$ is freely generated over $A(0)$ by the variables
    $x_{1}^{-i},x_{2}^{-i},\dots,x_{m_{i}}^{-i}\text{ in degree }-i\text{ for }i=1,2,\dots,2\ell,$
    $z_{1}^{-2\ell-1},z_{2}^{-2\ell-1},\dots,z_{m_{2\ell+1}}^{-2\ell-1}\text{ in degree }-2\ell-1,$
    $y_{1}^{k+i},y_{2}^{k+i},\dots,y_{m_{i}}^{k+i}\text{ in degree }k+i\text{ for }i=0,1,\dots,2\ell.$ (2.23)
    We then define an element $\omega^{0}=\sum_{i=0}^{2\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}+\sum_{j=1}^{m_{2\ell+1}}d_{dR}z_{j}^{-2\ell-1}d_{dR}z_{j}^{-2\ell-1}$ in $(\Lambda^{2}\Omega^{1}_{A})^{k}$, and set $\omega:=(\omega^{0},0,0,\dots)$ as before. Choose an element $H\in A^{k+1}$, the Hamiltonian, satisfying the analogue of classical master equation
    $\sum_{i=1}^{2\ell}\sum_{j=1}^{m_{i}}\dfrac{\partial H}{\partial x_{j}^{-i}}\dfrac{\partial H}{\partial y_{j}^{k+i}}+\frac{1}{4}\sum_{j=1}^{m_{2\ell+1}}\Big{(}\dfrac{\partial H}{\partial z_{j}^{-2\ell-1}}\Big{)}^{2}=0\text{ in }A^{k+2}.$ (2.24)
    Define $d$ on $A$ as in Equation( 2.21) with extra data $dz_{j}^{-2\ell-1}:=\dfrac{1}{2}\dfrac{\partial H}{\partial z_{j}^{-2\ell-1}}.$
    Finally, we define the element $\phi\in\Omega^{1}_{A})^{k}$ by
    $\phi$ $=\sum_{i=0}^{2\ell}\sum_{j=1}^{m_{i}}\big{[}-ix_{j}^{-i}d_{dR}y_{j}^{k+i}+(-1)^{i+1}(k+i)y_{j}^{k+i}d_{dR}x_{j}^{i}\big{]}$
    $+k\sum_{j=1}^{m_{2\ell+1}}z_{j}^{-2\ell-1}d_{dR}z_{j}^{-2\ell-1}.$ (2.25)
Observation 2.27.
In either case, the virtual dimension ${\rm vdim\ }A$ is always even. In fact, for any $k<0$ we have
${\rm vdim\ }A=\begin{cases*}2\sum_{i}(-1)^{i}m_{i},&$k$ \mbox{ even},\\ 0,&$k$ \mbox{ odd}.\end{cases*}$
Remark 2.28.
The classical Darboux theorem states that for a symplectic manifold $(X,\omega)$, one can find a local coordinate chart $(U;x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ such that $\omega|_{U}=\sum_{j}d_{dR}x_{j}d_{dR}y_{j}$. Moreover, we can write $\lambda=\sum_{j}x_{j}d_{dR}y_{j}$, called the Liouville form, such that $\omega|_{U}=d_{dR}\lambda$.
In this derived framework, the element $\phi$ above may seem to play the role of $\lambda.$ However, it is important to notice that $\phi$ is not a 1-form (of degree $k$) in the sense of Definition 2.17, because $d\phi\neq 0.$ Therefore, one needs to modify $\phi$ to obtain a genuine 1-form of degree $k$.
3 Shifted contact structures and a Darboux-type theorem
3.1 Basics of classical contact geometry
It is very well-known that contact manifolds are viewed as the odd-dimensional analogues of symplectic manifolds. In that respect, they have a number of common features: there is a Darboux theorem providing a local model for such structures; there is no local invariants; and it is more interesting to study their global properties. For more details, we refer to [17].
In this section, we shall revisit the basic aspects of contact geometry. There are in fact equivalent ways of describing the notion of a contact structure. We prefer to use the one below.
Definition 3.1.
Let $X$ be a manifold of dimension $12n+$ 1. A contact structure is a smooth field of tangent hyperplanes $\xi\subset TM$ (of rank ${\rm rank\ }2n$) with the property that for any smooth locally defining 1–form $\alpha$, i.e. $\xi=\ker(\alpha)$, the 2-form $d_{dR}\alpha|_{\xi}$ is non-degenerate.
Remark 3.2.
It is also possible to write $\xi=\ker(\alpha)$ with $\alpha$ a globally defined contact form on $M$ if and only if $\xi$ is coorientable, by which we mean that the quotient line bundle $TM/\xi$ is trivial. Except some pathological cases, it suffices to work with coorientable contact structures. More details can be found in [17, §1 & 2 ].
Note that if $d_{dR}\alpha|_{\xi}$ is non-degenerate, then for each $p\in X,$ $\xi_{p}=\ker(\alpha_{p})$ is a symplectic vector space with a symplectic form $\omega_{p}:=d_{dR}\alpha_{p}|_{\xi_{p}}.$ Therefore, we also call such 2-form $d_{dR}\alpha|_{\xi}$ symplectic.
It follows from the theory of symplectic vector spaces that $\dim\xi_{p}$ is even and the symplectic form on $\xi_{p}$ has a canonical form. It means that there exists a symplectic basis $\{e_{1},f_{1},\dots,e_{n},f_{n}\}$ for $\xi_{p}$ (and the corresponding dual basis $\{e_{1}^{*},f_{1}^{*},\dots,e_{n}^{*},f_{n}^{*}\}$ for $\xi_{p}^{*}$) satisfying
$\omega_{p}(e_{i},e_{j})=0=\omega_{p}(f_{i},f_{j})\text{ and }\omega_{p}(e_{i},f_{j})=-\omega_{p}(f_{j},e_{i})=\delta_{ij}\ \ \forall i,j,$ (3.1)
so that $\omega_{p}$ has the form $\omega_{p}=\sum_{i}e_{i}^{*}\wedge f_{i}^{*}.$
Let $(X,\xi=\ker(\alpha))$ be a contact manifold of dim $2n+1,$ and $p\in X$. Then we have a splitting
$T_{p}X=\xi_{p}\oplus\ker d_{dR}\alpha_{p}|_{\xi_{p}},$ (3.2)
where $\dim\xi_{p}=2n$ and $\dim\ker d_{dR}\alpha_{p}|_{\xi_{p}}=1$. In fact, as $d_{dR}\alpha|_{\xi}$ is non-degenerate, one can find a local trivialization $\{e_{1},f_{1},\dots,e_{n},f_{n},r\}$ of $TM=\ker\alpha\oplus rest$ such that
$\ker\alpha={\rm Span}\{e_{1},f_{1},\dots,e_{n},f_{n}\}\text{ and }rest={\rm Span}\{r\}.$ (3.3)
Moreover, using this splitting, one can find a unique vector field $\bf R$, called the Reeb vector field of $\alpha$, satisfying $\iota_{\bf R}d_{dR}\alpha=0$ and $\iota_{\bf R}\alpha=1.$
Example 3.3.
On $\mathbb{R}^{2n+1}$ with cartesian coordinates $(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n},z)$, the so-called standard contact form is given by
$\alpha_{std}=\displaystyle-d_{dR}z+\sum_{i=1}^{n}y_{i}d_{dR}x_{i}.$ (3.4)
Let $\xi\subset T\mathbb{R}^{2n+1}$ be the hyperplane field of rank ${\rm rank\ }2n$ defined by $\alpha_{std},$ i.e. $\xi=\ker\alpha_{std}.$ Then we observe that
$\ker(\alpha_{std})={\rm Span}\Big{\{}\dfrac{\partial}{\partial y_{j}},y_{j}\dfrac{\partial}{\partial z}+\dfrac{\partial}{\partial x_{j}}:j=1,\cdots,n\Big{\}}\ \text{ and }\ d_{dR}\alpha_{std}=\sum_{i}d_{dR}x_{i}\wedge d_{dR}y_{i}.$
Write $A_{j}={\partial/\partial y_{j}}$ and $B_{j}=y_{j}{\partial/\partial z}+{\partial/\partial x_{j}},$ then it is enough to observe that $d_{dR}\alpha_{std}(A_{i},B_{j})=\delta_{ij}$ and $d_{dR}\alpha_{std}(A_{i},A_{j})=0=d_{dR}\alpha_{std}(B_{i},B_{j}).$ It follows that $d_{dR}\alpha_{std}|_{\xi}$ is non-degenerate, and hence $\alpha_{std}$ is a contact form. Moreover, the Reeb vector field of $\alpha_{std}$ is $\textbf{R}={\partial/\partial z}.$
As in the symplectic case, there is a Darboux-type theorem for contact structures. It basically says that all contact structures can be locally given as in Equation (3.4). More formally, we have
Theorem 3.4 (Darboux Theorem for contact structures).
Let $(X,\alpha)$ be a contact manifold of dimension $12n+$ 1, and $p\in X$. Then there exists a local coordinate system $(U;x_{1},\cdots,x_{n},y_{1},\cdots,y_{n},z)$ around $p$ such that $p=(0,0,\cdots,0)$ and
$\alpha|_{U}=\displaystyle-d_{dR}z+\sum_{i=1}^{n}y_{i}d_{dR}x_{i}.$ (3.5)
3.2 Shifted contact structures and Darboux models
In this section, we provide an appropriate analogue of Definition 3.1 for derived spaces and study the local theory of shifted contact structures. Let $\bf X$ be a locally finitely presented derived (Artin) stack. Then we have:
Definition 3.5.
A pre-$k$-shifted contact structure on $\bf X$ consists of a perfect complex $\mathcal{K}$ on $\bf X$ with a morphism $\kappa:\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}$ of perfect complexes whose cone $Cone(\kappa)$ is of the form $L[k]$, up to quasi-isomorphism, where $L$ is a line bundle [^2] In the spirit of Remark 3.2, we say that a pre-$k$-shifted contact structure is coorientable if $L$ in the data is trivial.. Denote such a structure on $\bf X$ by $(\mathcal{K},\kappa,L)$.
Definition 3.6.
We say that a pre-$k$-shifted contact structure $(\mathcal{K},\kappa,L)$ on $\bf X$ is a $k$-shifted contact structure if locally on $\bf X$, where $L$ is trivial, the induced $k$-shifted 1-form $\alpha:\mathbb{T}_{X}\rightarrow\mathcal{O}_{X}[k]$ is such that the induced map $d_{dR}\alpha|_{\mathcal{K}}\ \cdot:=\kappa^{\vee}[k]\circ(d_{dR}\alpha\ \cdot)\circ\kappa:\mathcal{K}\rightarrow\mathcal{K}^{\vee}[k]$ is a weak equivalence. In that situation, we say the $k$-shifted 2-form $d_{dR}\alpha$ is non-degenerate on $\mathcal{K}$.
Moreover, when $k\leq 0$, the fiber-cofiber sequence $\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}\rightarrow L[k]$ in $QCoh(X)$ splits Zariski locally for any derived scheme. Call such local form a $k$-contact form.
Let $\bf X$ be a locally finitely presented derived Artin stack with a $k$-shifted contact structure $(\mathcal{K},\kappa,L)$. Recall from Yoneda’s lemma, ${\bf X}(A)\simeq Map_{dPstk}({\rm Spec}A,{\bf X})$, and hence any $A$-point $p\in{\bf X}(A)$ can be seen as a morphism $p:{\rm Spec}A\rightarrow\bf X$ of derived pre-stacks. Then, let us consider the pair $(p,\alpha_{p})$, with $p\in{\bf X}(A),\ \alpha_{p}\in p^{*}(\mathbb{L}_{{\bf X}}[k])$ a $k$-contact form. For $A\in cdga_{{\mathbb{K}}}$, there is a $\mathbb{G}_{m}(A)$-action on the pair $(p,\alpha_{p})$ by
$f\triangleleft(p,\alpha_{p}):=(p,f\cdot\alpha_{p}).$
Denote by $H^{0}$ the functor sending $A\mapsto H^{0}(A).$ Denote the image under $H^{0}$ of an element $f$ simply by $f^{0}.$ Note that localizing $A$ if necessary, w.l.o.g. we may assume that the image $f^{0}$ is always invertible. It follows that $f^{0}$ lies in $(A^{0})^{\times}$, which is by definition $\mathbb{G}_{m}(A^{0})=(A^{0})^{\times}$.
Observation 3.7.
If ${\bf X},(p,\alpha_{p})$, and the $\mathbb{G}_{m}(A)$-action are as above, then for an element $f\in\mathbb{G}_{m}(A)$, we can obtain $Cocone(f\cdot\alpha_{p})\simeq Cocone(\alpha_{p})$ by using the invertibility of $f.$
Because our results in this paper are all about the local structure theory of contact derived stacks, we will focus on the refined affine case in the sense of Theorem 2.5. In this regard, we will always assume that our refined local models are given in terms of minimal standard form cdgas.
From Proposition 2.15, on a refined affine neighborhood, say on ${\rm Spec}A$ with $A$ a minimal standard form cdga, the perfect complexes $\mathbb{T}_{A},\mathbb{L}_{A}$, when restricted to ${{\rm Spec}H^{0}(A)}$, are both free finite complexes of $H^{0}(A)$-modules. In that case, Definitions 3.5 and 3.6, and Observation 3.7 will reduce to the following local descriptions, where $\mathcal{K},Cone(\kappa)$ will be equivalent to the ordinary $\ker(\alpha),{\rm coker}(\kappa)$, respectively in $D(Mod_{A})$; and $L$ in the splitting corresponds to the line bundle generated by the Reeb vector field of the classical case.
Shifted contact structures with (nice) local models.
Recall that the (mapping) cone of a morphism $f:A\rightarrow B$ in some homotopical category is a realization of the homotopy cofiber of $f$. That is, it is the homotopy pushout satisfying universally the homotopy diagram
[…]
It is also called the homotopy cokernel of $f$ or the weak quotient of $B$ by the image of $A$ under $f$.
As a dual notion, the (mapping) cocone of a morphism $f:A\rightarrow B$, with $B$ a pointed object, in a homotopical category is a particular realization of the homotopy fiber of $f$ (i.e. of the homotopy pullback of the point along $f$). In that case, the following diagram homotopy commutes:
[…]
It is also called the homotopy kernel of $f$.
Remark 3.8.
In the case of a morphism (perfect) of complexes $f:A\rightarrow B$, we have =0$\star$=0⋆ = 0 (as the initial and terminal objects) and that the complex $\ker(f)$, which is the subcomplex of $A$ formed by the (strict) kernels $\{\ker(f_{n})\},$ commutes the diagram
[…]
which essentially means $f\circ i=0$. This is in fact a strict pullback diagram. From the universality of the homotopy kernel of $f$, there is a natural inclusion $\ker(f)\hookrightarrow Cocone(f)$ from the strict fiber to the homotopy fiber (i.e. from the strict kernel to the homotopy kernel). This map is in fact an equivalence if $f$ is a fibration (i.e. a surjective morphism of complexes).
${\rm coker}(f)$, on the other hand, is the quotient complex of $B$ formed by the cokernels $\{{\rm coker}(f_{n})\}.$ By definition, ${\rm coker}(f)$ commutes the diagram
[…]
From the universality of the homotopy cokernel of $f$, there is a natural map $Cone(f)\rightarrow{\rm coker}(f)$ from the homotopy cofiber to the strict cofiber (i.e. from the homotopy cokernel to the strict cokernel). This map is in fact an equivalence if $f$ is a cofibration (i.e. an injective morphism of complexes).
Now, in the presence of Remark 3.8, we revisit Definitions 3.5 and 3.6 with (nice) local models. Let $\textbf{X}={\rm Spec}A$ be an affine derived $\mathbb{K}$-scheme for $A$ a minimal standard form cdga, endowed with a $k$-shifted contact structure $(\mathcal{K},\kappa,L)$ for $k<0$. Assume also that $L$ is trivial on ${\rm Spec}A.$ From definitions, the triangle $\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}\rightarrow L[k]$ splits and $Cone(\kappa)$ is of the form a line bundle $L[k]$. In that case, we have the homotopy commuting square
[…]
If $\kappa:\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}$ is a monomorphism, the homotopy cofibers are equivalent to strict cofibers by Remark 3.8. Thus, $Cone(\kappa)$ is equivalent to the ordinary ${\rm coker}(\kappa)$.
On $\textbf{X}={\rm Spec}A$, where $L$ is trivial, $L[k]$ is of the form $\mathcal{O}[k].$ Recall that since $A$ is a minimal standard form cdga, we have $\mathbb{L}_{{\bf X}}\simeq\mathbb{L}_{A}$ in $D(Mod_{A})$ and the perfect complexes $\mathbb{T}_{A},\mathbb{L}_{A}$, when restricted to ${{\rm Spec}H^{0}(A)}$, are both free finite complexes of $H^{0}(A)$-modules. Let $\alpha$ be a $k$-contact form, with the underlying $k$-shifted 1-form $\alpha:\mathbb{T}_{A}\rightarrow A[k]$, then we have
[…]
Here both outer and inner squares (resp. the homotopy fiber and the homotopy cofiber) commute, and $\mathcal{K}$ is equivalent to $Cocone(\underline{\alpha})$, with $\underline{\alpha}\in\mathcal{A}^{1}({\rm Spec}A,k)$ a map [^3] The map from the epi-mono factorization of $\alpha:\mathbb{T}_{A}\twoheadrightarrow{\rm Im}\alpha\hookrightarrow\mathcal{O}[k]$, where $\underline{\alpha}:\mathbb{T}_{A}\rightarrow{\rm Im}\alpha.$ of perfect complexes. From Remark 3.8, there is a natural map $\ker(\underline{\alpha})\hookrightarrow\mathcal{K}$ from the strict kernel to the homotopy kernel of $\underline{\alpha}$. Since $\underline{\alpha}$ is an epimorphism, the natural map $\ker(\underline{\alpha})\hookrightarrow\mathcal{K}$ is then an equivalence. Thus, on suitable local models, we can use the strict kernel to represent the homotopy kernel.
Now, we consider an explicit general form of the underlying $k$-shifted 1-form $\alpha:\mathbb{T}_{A}\rightarrow A[k]$ on ${\rm Spec}A$ with $A$ as above. Note that it is the minimal (at $p\in{\rm Spec}H^{0}(A)$) compared to all other cdgas quasi-isomorphic to $A$ (at $p\in{\rm Spec}H^{0}(A)$). Explicitly, Letting $A=A(-k)$, $A$ is the free graded algebra over $A(0)$ generated by the variables $x_{1}^{-i},x_{2}^{-i},\dots,x_{m_{i}}^{-i}$, with $m_{i}\in\mathbb{Z}$, for $i=1,\dots,-k$ such that $d_{dR}x_{1}^{-i},d_{dR}x_{2}^{-i},\dots,d_{dR}x_{m_{i}}^{-i},\ i=1,\dots,-k,$ is a $A$-basis for $\Omega_{A}^{1}$. Write $\underline{\alpha}=\alpha=\sum_{i,j}\alpha_{j}^{k+i}d_{dR}x_{j}^{-i}$ with $\alpha_{j}^{n}\in A^{n}$. Notice that by definition, we have $\ker(\alpha)=\ker(\underline{\alpha})$. Thus, when we consider the strict kernel on such local models, we simply write $\ker(\alpha).$
Therefore, the discussion above implies that the data of (1) a submodule $\mathcal{K}$ of $\mathbb{T}_{A}=Der(A)$ with the natural inclusion $i:\mathcal{K}\hookrightarrow Der(A)=(\Omega^{1}_{A})^{\vee}$; (2) a morhism $\mathbb{T}_{A}\rightarrow L[k]\simeq{\rm coker}(i)$ for $L[k]$ a shifted line bundle; and (3) the non-degeneracy condition on the induced map $\mathbb{T}_{A}\rightarrow\mathcal{O}[k]$ will define a shifted contact structure on $\textbf{X}={\rm Spec}A$, where $A$ is a (minimal) standard form cdga. That is, we obtain the following proposition/definition for nice affine derived schemes:
Definition with nice local models. For a (minimal) standard form cdga $A$ and $k<0$, a $k$-shifted (strict) contact structure on ${\rm Spec}A$ is a submodule $\mathcal{K}$ of $\mathbb{T}_{A}=Der(A)$ with the natural inclusion $i:\mathcal{K}\hookrightarrow Der(A)=(\Omega^{1}_{A})^{\vee}$ such that $\mathcal{K}\simeq\ker(\alpha)$ for a $k$-shifted 1-form $\alpha$ with the property that the $k$-shifted 2-form $d_{dR}\alpha$ is non-degenerate on $\ker(\alpha)$ and the complex $Cone(i:\mathcal{K}\hookrightarrow\mathbb{T}_{A})\simeq{\rm coker}(i)$ is the quotient complex and of the form $L[k]$, with $L$ a line bundle.
Observation 3.9.
Let $\textbf{X}={\rm Spec}A$ be an affine derived $\mathbb{K}$-scheme for $A$ a minimal standard form cdga. For any $f\neq 0$ in $A$ and any $k$-shifted contact form $\alpha$, one has $\ker(\alpha)\simeq\ker(f\alpha).$ Hence, both define equivalent contact structures on $\bf X$. In fact, this follows from the fact that the contraction operation $\iota_{Y}$ on the de Rham algebra $DR(A)$ with a homogeneous vector field $Y$ is the unique derivation of degree $|Y|+1$ such that $\iota_{Y}g=0$ and $\iota_{Y}d_{dR}g=Y(g)$ for all $g\in A.$ Therefore,
$\iota_{Y}(f\alpha)=(\iota_{Y}f)\cdot\alpha+(-1)^{|Y|+1}f\cdot\iota_{Y}\alpha=(-1)^{|Y|+1}f\cdot\iota_{Y}\alpha.$
Adopting the classical terminology (in terms of non-integrable distributions), on refined local models, we sometimes call the subcomplex $\ker(\alpha)$ of $\mathbb{T}_{{\bf X}}$ a contact structure and the corresponding $k$-shifted 1-form $\alpha$ a (locally) defining $k$-contact form.
In what follows, we give a prototype construction for $k$-shifted contact forms, which is similar to the previous case of shifted symplectic Darboux models. In brief, we will coherently extend the symplectic case studied in [7, Example 5.8] using similar constructions.
Example 3.10.
In this example, fixing $\ell\in{\mathbb{N}}$, we construct a standard form cdga $A=A(n)$ with the induced $k$-shifted contact structure on ${\rm Spec}A$, for $k=-2\ell-1=-n$.
Step-1: Construction of a cdga $(A,d)$. First, we consider a smooth ${\mathbb{K}}$-algebra $A(0)$ of dimension $m_{0}$. We assume that there exist degree 0 variables $x_{1}^{0},x_{2}^{0},\dots,x_{m_{0}}^{0}$ in $A(0)$ such that $d_{dR}x_{1}^{0},\dots,d_{dR}x_{m_{0}}^{0}$ form a basis for $\Omega_{A(0)}^{1}$ over $A(0).$ This choice can be made by localizing $A(0)$ if necessary. Next, choosing non-negative integers $m_{1},\dots,m_{\ell}$, define a commutative graded algebra $A$ to be the free graded ${\mathbb{K}}$-algebra over $A(0)$ generated by the variables
[…]
where we call $z^{k}$ the distinguished variable (of $\deg k$). It follows that $\Omega^{1}_{A}$ is the free $A$-module of finite rank with an $A$-basis $\big{\{}d_{dR}x_{j}^{-i},d_{dR}y_{j}^{k+i},d_{dR}z^{k}:i=0,1,\dots,\ell,\ j=1,2,\dots,m_{i}\big{\}}.$ Choose an element $H\in A^{k+1}$, called the Hamiltonian, satisfying the condition (the classical master equation)
$\sum_{i=1}^{\ell}\sum_{j=1}^{m_{i}}\dfrac{\partial H}{\partial x_{j}^{-i}}\dfrac{\partial H}{\partial y_{j}^{k+i}}=0\text{ in }A^{k+2}.$
Then we define the internal differential on $A$ by the equations $d|_{A(0)}=0;\ dx_{j}^{-i}={\partial H}/{\partial y_{j}^{k+i}}$ and $dy_{j}^{k+i}={\partial H}/{\partial x_{j}^{-i}}$ for all $i,j$; and $dz^{k}=0$. As discussed before, the condition on $H$ above is equivalent to saying ``$dH=0"$. Here, due to degree reasons, $H$ does not involve any of $z^{k},y^{k}_{j}$’s.
Notice that, by construction, $(A,d)$ is a standard form cdga with $A=A(n=2\ell+1)$ which is defined inductively by adjoining free modules $M^{-i}=\langle x_{1}^{-i},x_{2}^{-i},\dots,x_{m_{i}}^{-i}\rangle_{A(i-1)}$ for $i=1,2,\dots,\ell$; $M^{k+i}=\langle y_{1}^{k+i},y_{2}^{k+i},\dots,y_{m_{i}}^{k+i}\rangle_{A(-k-i-1)}$ for $i=1,\dots,\ell;$ and $M^{k}=\langle y_{1}^{k},y_{2}^{k},\dots,y_{m_{0}}^{k},z^{k}\rangle_{A(-k-1)}$. Moreover, the cdga $A$ has a virtual dimension ${\rm vdim\ }(A)=-1$.
Step-2: Pre-contact data. Define an element $\alpha\in(\Omega_{A}^{1})^{k},$ called the $k$-contact element, by
$\alpha=k\Bigg{[}d_{dR}z^{k}+\displaystyle\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}x_{j}^{-i}d_{dR}y_{j}^{k+i}\Bigg{]}.$ (3.13)
Then we get $d_{dR}\alpha=\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}\in(\Lambda^{2}\Omega^{1}_{A})[k]$. Let us first show that there is a natural pre-$k$-shifted contact structure on ${\rm Spec}A$ defined by $\ker\alpha.$ Denote the vector fields annihilating $\alpha$ by
$\zeta_{j}^{-i}$ $=\partial/\partial x_{j}^{-i}$ for $\text{for }i=0,\dots,\ell,\text{ and }1\leq j\leq m_{i},$
$\eta^{k+i}_{j}$ $=\partial/\partial y^{k+i}_{j}-x^{-i}_{j}\partial/\partial z^{k}$ for $\text{for }i=0,\dots,\ell,\text{ and }1\leq j\leq m_{i}.$ (3.14)
[…]
$\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}=\ker\alpha|_{{\rm spec}H^{0}(A)}\oplus Rest|_{{\rm spec}H^{0}(A)},$ (3.15)
where $\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}$ is a complex of free $H^{0}(A)$-modules by Proposition 2.15. Hence, we get a pre-$k$-shifted contact structure on ${\rm Spec}A$ in the sense of Definition 3.5, where for each homogeneous degree $n$, we have $(\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)})^{n}=(\ker\alpha|_{{\rm spec}H^{0}(A)})^{n}\oplus Rest^{n}|_{{\rm spec}H^{0}(A)}$ so that
$\ker\alpha|_{{\rm spec}H^{0}(A)}$ $=\langle\zeta_{j}^{-i},\ \eta^{k+i}_{j}:\ i=0,1,\dots\ell,\ j=1,\dots,m_{i}\rangle_{A(0)},$
$Rest|_{{\rm spec}H^{0}(A)}$ $=\langle\partial/\partial z^{k}\rangle_{A(0)}.$
Step-3: From pre-contact to contact. It remains to show that $d_{dR}\alpha$ is non-degenerate on ${\ker\alpha}$. To this end, it suffices to prove the non-degeneracy of the map $d_{dR}\alpha|_{\ker\alpha}\otimes_{A}\mathrm{id}_{H^{0}(A)}:\ker\alpha\otimes_{A}H^{0}(A)\rightarrow(\ker\alpha)^{\vee}[k]\otimes_{A}H^{0}(A).$ We first observe that, at $p\in{\rm spec}H^{0}(A)$, $d_{dR}\alpha|_{p}$ maps
$\langle\partial/\partial x_{1}^{-i}|_{p},\dots,\partial/\partial x_{m_{i}}^{-i}|_{p}\rangle_{{\mathbb{K}}}\xrightarrow{\sim}\langle d_{dR}y_{1}^{k+i}|_{p},\dots,d_{dR}y_{m_{i}}^{k+i}|_{p}\rangle_{{\mathbb{K}}},$
$\langle\partial/\partial y^{k+i}_{1}|_{p},\dots,\partial/\partial y^{k+i}_{m_{i}}|_{p}\rangle_{{\mathbb{K}}}\xrightarrow{\sim}\langle d_{dR}x_{1}^{-i}|_{p},\dots,d_{dR}x_{m_{i}}^{-i}|_{p}\rangle_{{\mathbb{K}}}$
isomorphically for all $i$, with $d_{dR}\alpha(-,\partial/\partial z^{k}|_{p})=0$. It follows that, at $p\in{\rm Spec}H^{0}(A)$, we have the identifications
[…]
$\simeq\langle\partial/\partial x_{j}^{-i}|_{p},\partial/\partial y^{k+i}_{j}|_{p}-(x^{-i}_{j}\partial/\partial z^{k})|_{p}:\forall i,j\rangle_{{\mathbb{K}}}\simeq(\ker\alpha)|_{p}.$ (3.16)
Thus, the maps $(d_{dR}\alpha|_{\ker\alpha})^{-i}|_{p}:(\ker\alpha|_{p})^{-i}\rightarrow(\ker^{\vee}\alpha|_{p})^{k+i}$ are isomorphisms at $p$, hence in a neighborhood of $p$. So, localizing $A$ at $p$ if necessary, the map $d_{dR}\alpha|_{\ker\alpha}\otimes_{A}\mathrm{id}_{H^{0}(A)}$ is an isomorphism of complexes, and hence a quasi-isomorphism. This proves the non-degeneracy. Thus, we get a $k$-shifted contact structure on ${\rm Spec}A$ in the sense of Definition 3.6 as desired.
Step-4: $k$-contact forms. From [7, Example 5.8],   $d_{dR}\alpha$ is closed w.r.t both $d$ and $d_{dR}$ such that there is an element $\phi\in(\Omega^{1}_{A})^{k}$ satisfying
$dH=0\text{ in }A^{k+2},\ d_{dR}H+d\phi=0\text{ in }(\Omega^{1}_{A})^{k+1},\text{ and }d_{dR}\phi=kd_{dR}\alpha.$ (3.17)
We can set $\phi:=k\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}x_{j}^{-i}d_{dR}y_{j}^{k+i}$ and write $\alpha=kd_{dR}z^{k}+\phi$. Notice that $\alpha\in\Omega^{1}_{A}[k]$ is a representative of $\phi$, but with $d\phi=-d_{dR}H$. Leaving $d_{dR}\alpha$ unchanged, we can actually make $\alpha$ $d$-closed via a minor modification: Choose an element $\theta\in A^{k}$ such that $d\theta=H$ in $A^{k+1}$ and write
$\alpha_{0}=-d_{dR}\theta+\alpha.$ (3.18)
Using (3.17), it is now straightforward to check that $\alpha_{0}$ is $d$-closed, and hence a 1-form of degree $k.$
Finally, using our coordinates, we can also write $d_{dR}\theta=\theta^{\prime}d_{dR}z^{k}+\sum\theta_{j}d_{dR}y_{j}^{k}$, with $\theta^{\prime},\theta_{j}\in A(0)$ due to degree reasons. Then $\alpha_{0}$ takes the form
$\alpha_{0}=c_{1}d_{dR}z^{k}+\displaystyle k\Bigg{[}\sum_{i=0,j=1}^{m_{0}}(c_{j}+x_{j}^{0})d_{dR}y_{j}^{k}+\sum_{i=1}^{\ell}\sum_{j=1}^{m_{i}}x_{j}^{-i}d_{dR}y_{j}^{k+i}\Bigg{]},$ (3.19)
where $c_{1}=-\theta^{\prime}+k,c_{j}=-\theta_{j}/k$. Now, $\ker\alpha_{0}$ is spanned by similar vector fields as above with obvious modifications and induces a $k$-shifted contact structure in a similar way.
Definition 3.11.
If $A,\alpha,\alpha_{0}$ with the variables $x_{j}^{-i},y_{j}^{k+i},z^{k}$ are as above, we then say $A,\alpha$ (or $\alpha_{0}$) are in contact Darboux form.
Remark 3.12.
Note that the expression in Equation (3.18) will still be valid for the other cases $(a)\ k\equiv 0\mod 4,\text{ and }(b)\ k\equiv 2\mod 4$. Equations ((a)) -- ((b)) show that the other cases in fact involve modified versions of $H,d,$ and $\phi$ with some possible extra terms. In any case, the modified $A,\alpha_{0}$ would also serve as the desired contact model. Following the same terminology as above, we would again say $A,\alpha_{0}$ are in (contact) Darboux form.
3.3 A Darboux-type theorem for negatively shifted contact derived schemes
In what follows, we give the proof of Theorem 1.1, which essentially says that for $k<0$, every $k$-shifted contact derived ${\mathbb{K}}$-scheme $\bf X$ is locally equivalent to $({\rm Spec}A,\alpha_{0})$ for $A$ a minimal standard form cdga and $\alpha_{0}$ as in Example 3.10 (with some minor modifications depending on $k$, if necessary). Our result is a “contact" variation of [7, Theorem 5.18.]. More precisely, we have:
Theorem 3.13.
Let $\bf X$ be a (locally finitely presented) derived $\mathbb{K}$-scheme with a $k$-shifted contact structure $(\mathcal{K},\kappa,L)$ for $k<0$, and $x\in\bf X$. Then there is a local contact model $\big{(}A,\alpha_{0}\big{)}$ and $p\in{\rm spec}H^{0}(A)$ such that $i:{\rm Spec}A\hookrightarrow\bf X$ is an open inclusion with $i(p)=x$,  $A$ is a standard form cdga that is minimal at $p$, and $\alpha_{0}$ is a $k$-shifted contact form on ${\rm Spec}A$ such that $A,\alpha_{0}$ are in standard contact Darboux form.
Note that for $k<0$ odd, the pair $(A,\alpha_{0})$ is explicitly described in Example 3.10, see Equations (3.10), (3.13), and (3.19). For the other cases, one should use another sets of variables as in Equations ((a)) and ((b)), and modify $H,\phi,d$ accordingly.
Before giving the proof, we begin by some remarks and simplifying assumptions.
Remark 3.14.
We first note that as $\bf X$ is a locally finitely presented derived $\mathbb{K}$-scheme, there exists a cover by affine derived $\mathbb{K}$-subschemes of finite presentation. Thus, for $x\in\bf X$ we can choose a cdga $B$ of finite presentation with a Zariski open inclusion $\iota:{\rm Spec}B\hookrightarrow\bf X$ and a unique $q\in{\rm spec}H^{0}(B)$ such that $q\mapsto x$. In this case, $B$ being of finite presentation implies that $\mathbb{L}_{B}$ has finite Tor-amplitude [^4] We say that a perfect complex $E$ of $R$-modules has Tor-amplitude in some interval $[a,b]$ if $H^{i}(E\otimes_{R}^{L}N)=0$ for all $i\notin[a,b]$ and for all $R$-modules $N$., say in $[-n,0]$ for some $n\in\mathbb{Z}_{+}$. Then it follows from [12, Prop. 7.2.4.23] that for any integer $k<0$,  $\mathbb{L}_{B}[-k-n]$ has Tor-amplitude in $[k,0]$. Notice that both complexes are equivalent. Since we will be interested in $k$-shifted structures, for the proof, we equivalently use this shifted complex of $B$. Therefore, while getting a refined neighborhood, we assume w.l.o.g. that for $k<0$, the corresponding cdga $B$ is such that its cotangent complex $\mathbb{L}_{B}$ has Tor-amplitude in $[k,0].$
Remark 3.15.
Let ${\bf X},x,B,q$ be as in Remark 3.14. Then the construction given in [7, Theorem 4.1] (cf. Theorem 2.5) ensures that there exists a suitable localization of $B$ at $q$, which is equivalent to a minimal standard form cdga $A=A(m)$, for some $m$, constructed inductively as in (2.2) such that there exists $p\in{\rm Spec}A$ with $p\mapsto q$. Here the integer $m$ is determined by the Tor-amplitude of $\mathbb{L}_{B}$, which is by assumption $\leq-k$ [^5] It means the perfect complex has Tor-amplitude in $[k,0]$. It should also be noted that during induction, each $\mathbb{L}_{A(\ell)}$ has Tor-amplitude in $[-\ell,0]$. Moreover, as there is an equivalence $A(-k)\rightarrow B$, Proposition 2.15 provides a simple description for $\mathbb{L}_{A}$, with Tor-amplitude $\leq-k$. That is, when restricted to ${\rm spec}H^{0}(A)$, $\mathbb{L}_{A}$ is equivalent to the complex $\Omega_{A}^{1}\otimes_{A}H^{0}(A)$ of free $H^{0}(A)$-modules
$0\rightarrow V^{k}\rightarrow V^{k+1}\rightarrow\cdots\rightarrow V^{-1}\rightarrow V^{0}\rightarrow 0,$
with $d|_{p}^{-i}=0$ for $i=1,2,\dots,-k$ (due to the minimality of $A$).
Proof of Theorem 3.13.
The proof is essentially based on that of [7, Theorem 5.18.]. Let $k<0$ and $x\in\bf X$, apply Theorem 2.5 to get a refined neighborhood ${\bf U}={\rm Spec}A$ of $x$ with $p\in{\rm spec}H^{0}(A)$ such that $i:{\rm Spec}A\hookrightarrow\bf X$ is an open inclusion, $i(p)=x$, and $A$ is a standard form cdga that is minimal at $p$. We also assume that $A$ is in fact constructed inductively as described in (2.2) with $A=A(-k)$ such that $\mathbb{L}_{A}$ has Tor-amplitude in $[k,0]$.
W.l.o.g., we also assume that $L$ is trivial on $\bf U$, and hence, over $\bf U$, the induced 1-form $\alpha:\mathbb{T}_{X}\rightarrow\mathcal{O}_{X}[k]$ is such that $\mathcal{K}$ is the cocone of ${\alpha}$, up to quasi-isomorphism, and the 2-form $d_{dR}\alpha$ is non-degenerate on $\mathcal{K}$. In that case, the triangle $\mathcal{K}\rightarrow\mathbb{T}_{{\bf X}}\rightarrow L[k]$ splits over $\bf U.$ That is, $\mathrm{cofib}(i^{*}\mathcal{K}\rightarrow i^{*}\mathbb{T}_{{\bf X}})\simeq i^{*}(L[k])$, which is concentrated in $\deg-k$, and $i^{*}(\mathbb{T}_{{\bf X}})=i^{*}(\mathcal{K})\oplus i^{*}(L[k]).$ Here, we will call the 2nd summand $Rest$. Note also that $i^{*}$ is just the restriction to ${\bf U}={\rm Spec}A$, so we will omit it in the notation or use $(-)|_{{\rm Spec}A}$ instead.
We fix the locally defining 1-form $\alpha$ over $\bf U$ for the rest of the proof. From now on, we will use the properties of shifted contact structures when restricted to nice local models. I.e., over ${\rm Spec}A$, we use the strict $\ker,{\rm coker}$ in $D(Mod_{A}).$
Consider the sequence $\omega_{u}:=(d_{dR}\alpha,0,0,\dots),$ which defines a closed $k$-shifted 2-form on ${\bf U}$ in the sense of Definition 2.20. Applying Proposition 2.24 to $k\omega_{u}$, we obtain elements $H\in A^{k+1}$ and $\phi\in(\Omega^{1}_{A})^{k}$ such that $dH=0$,  $d_{dR}H+d\phi=0$ , and $k\omega_{u}\sim(d_{dR}\phi,0,0,\dots).$ Notice that we in fact have $d_{dR}\phi=kd_{dR}\alpha,$ because there is no non-trivial $\beta\in(\Lambda^{2}\Omega^{1}_{A})^{k-1}$ satisfying the relation $kd_{dR}\alpha-d_{dR}\phi=d\beta$ due to degree reasons.
From Proposition 2.15, the tangent complex $\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}=(\mathbb{L}_{A}|_{{\rm spec}H^{0}(A)})^{*}$ is also represented by a complex of free finite rank $H^{0}(A)$-modules, with Tor-amplitude in $[0,-k]$. Then the $k$-shifted 2-form $d_{dR}\alpha$ defines the induced map of perfect complexes via $v\mapsto\iota_{v}d_{dR}\alpha$:
[…]
where both horizontal differentials $d^{i},(d^{i})^{*}$ are zero at $p\in{\rm spec}H^{0}(A)$ due to the minimality of $A$.
From the contactness condition, $d_{dR}\alpha$ is non-degenerate on the subcomplex $\ker\alpha|_{{\rm spec}H^{0}(A)}$ of $\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}$. Therefore, the quotient ${\mathbb{T}_{A}}/{\ker\alpha}$ can also be identified with a $k$-shifted line bundle $Rest$ (concentrated in degree $-k$) such that one has the splitting
$\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}=\ker\alpha|_{{\rm spec}H^{0}(A)}\oplus Rest|_{{\rm spec}H^{0}(A)}.$ (3.21)
Write $W$ for the dual of $\ker\alpha|_{{\rm spec}H^{0}(A)}$ in $\mathbb{L}_{A}|_{{\rm spec}H^{0}(A)}$, i.e. $W:=(\ker\alpha|_{{\rm spec}H^{0}(A)})^{*}$, then we have the diagram
[…]
such that the vertical maps $d_{dR}\alpha|_{\ker\alpha}:(W^{k+i})^{*}\rightarrow W^{-i}$, $v\mapsto\iota_{v}d_{dR}\alpha$, are all quasi-isomorphisms.
Observation 3.16.
As both horizontal differentials $d^{i},(d^{i})^{*}$ are zero at $p\in{\rm spec}H^{0}(A)$ because of the minimality of $A$, the vertical maps are isomorphisms at $p$, and hence isomorphisms in a neighborhood of $p$. By localizing $A$ at $p$ if needed, we may assume that the vertical maps are all isomorphisms.
When $k$ is odd.
We now focus on a particular and the simplest case: $k$ is odd. Let $k=-2\ell-1$ for $\ell\in{\mathbb{N}}$. Localizing $A$ at $p$ if necessary, first choose degree 0 variables $x_{1}^{0},x_{2}^{0},\dots,x_{m_{0}}^{0}$ in $A(0)$ such that $\{d_{dR}x_{j}^{0}:j=1,\dots,m_{0}\}$ forms a basis for $W^{0}$ over $A(0),$ and $(Rest^{*})^{0}=0$
Now, by the equivalence $d_{dR}\alpha:(W^{k+i})^{*}\rightarrow W^{-i}$, we have $H^{-i}(W)\simeq H^{k+i}(W)^{*},$ and hence $\dim H^{-i}(W)=\dim H^{k+i}(W).$ Then we can choose the following set of generators for the free algebra $A$:
  • When $i=0$, we can find generators $y_{1}^{k},y_{2}^{k},\dots,y_{m_{0}}^{k},z^{k}\in A^{k}$ such that $\{d_{dR}y_{1}^{k},\dots,d_{dR}y_{m_{0}}^{k}\}$ is a basis for $W^{k}$ which is dual to the basis $\{d_{dR}x_{1}^{0},\dots,d_{dR}x_{m_{0}}^{0}\}$ for $W^{0}$ and that the complex $Rest$, which is concentrated in $\deg-k$, is generated by the vector field $\partial/\partial z^{k}$ of degree $-k$.
  • Similarly, for $1\leq i\leq\ell$, choose $x_{1}^{-i},\dots,x^{-i}_{m_{i}}\in A^{-i};\text{ and }y_{1}^{k+i},\dots,y_{m_{i}}^{k+i}\in A^{k+i}$ such that $\{d_{dR}y_{j}^{k+i}:j=1,2,\cdots,m_{i}\}$ is a basis for $W^{k+i}$ over $A(0)$ which is dual to the basis $\{d_{dR}x_{1}^{-i},\dots,d_{dR}x_{m_{i}}^{-i}\}$ for $W^{-i}$.
That is, using the local coordinates above, for $\text{for }i=0,1,\dots,\ell,$ we have
$W^{-i}$ $=\langle d_{dR}x^{-i}_{1},d_{dR}x^{-i}_{2},\dots,d_{dR}x^{-i}_{m_{i}}\rangle_{H^{0}(A)},$
$W^{k+i}$ $=\langle d_{dR}y^{k+i}_{1},d_{dR}y^{k+i}_{2},\dots,d_{dR}y^{k+i}_{m_{i}}\rangle_{H^{0}(A)}.$
By Observation 3.16, the isomorphisms in Diagram (3.22) imply that for $i=0,1,\dots,\ell,$ we have
$(W^{-i})^{*}$ $=\langle\partial/\partial x_{1}^{-i},\dots,\partial/\partial x_{m_{i}}^{-i}\rangle_{H^{0}(A)}\xrightarrow{\sim}\langle d_{dR}y^{k+i}_{1},\dots,d_{dR}y^{k+i}_{m_{i}}\rangle_{H^{0}(A)},$ (3.23)
$(W^{k+i})^{*}$ $=\langle\partial/\partial y^{k+i}_{1},\dots,\partial/\partial y^{k+i}_{m_{i}}\rangle_{H^{0}(A)}\xrightarrow{\sim}\langle d_{dR}x_{1}^{-i},\dots,d_{dR}x^{-i}_{m_{i}}\rangle_{H^{0}(A)}.$ (3.24)
Then, by the splitting in Equation (3.21), we have
$\ker\alpha|_{{\rm spec}H^{0}(A)}$ $=\big{\langle}\partial/\partial x_{1}^{-i},\dots,\partial/\partial x_{m_{i}}^{-i},\partial/\partial y^{k+i}_{1},\dots,\partial/\partial y^{k+i}_{m_{i}}:i=0,1,\dots,\ell\ \big{\rangle}_{H^{0}(A)},$
$Rest|_{{\rm spec}H^{0}(A)}$ $=\big{\langle}\partial/\partial z^{k}\big{\rangle}_{H^{0}(A)}.$
Here $Rest|_{{\rm spec}H^{0}(A)}$ is a subcomplex of $\mathbb{T}_{A}|_{{\rm spec}H^{0}(A)}$ that is concentrated in degree $-k$ and generated by the degree $-k$ vector field $\partial/\partial z^{k}$. Moreover, $\ker\alpha$ has $\mathrm{Tor}$-amplitude $[0,-k]$.
$A$ is now identified with the standard form cdga over $A(0)$ freely generated by the variables $z^{k},x_{j}^{-i},y_{j}^{k+i}$ as in Example 3.10. We also impose suitable differential $d$ as before: $d$ acts on $x_{j}^{-i},y_{j}^{k+i}$ as in Equation (2.21) such that we additionally set $dz^{k}=0$.
The non-degeneracy condition on $d_{dR}\alpha|_{\ker\alpha}$ sending dual basis of $d_{dR}x_{1}^{-i},\dots,d_{dR}x^{-i}_{m_{i}}$ to the basis $d_{dR}y^{k+i}_{1},\dots,d_{dR}y^{k+i}_{m_{i}}$ (and vice versa) as in Equations (3.23) and (3.24) implies that
$d_{dR}\alpha|_{\ker\alpha}=\displaystyle\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}.$ (3.25)
[…]
$(\ker\alpha)^{k+i}|_{{\rm spec}H^{0}(A)}$ $=\langle\partial/\partial y^{k+i}_{1},\dots,\partial/\partial y^{k+i}_{m_{i}}\rangle_{H^{0}(A)}$
$\simeq\langle\partial/\partial y^{k+i}_{1}-x^{-i}_{1}\partial/\partial z^{k},\dots,\partial/\partial y^{k+i}_{m_{i}}-x^{-i}_{m_{i}}\partial/\partial z^{k}\rangle_{H^{0}(A)},$
where $(\ker\alpha)^{-i}|_{{\rm spec}H^{0}(A)}=\langle\partial/\partial x_{j}^{-i}:1\leq j\leq m_{i}\rangle_{H^{0}(A)}.$ On the other hand, [7, Example 5.8] shows that we may explicitly write [^6] More generally, using the defining equations in (3.17), one can write $\phi=\sum_{i,j}(-i)x_{j}^{-i}d_{dR}y_{j}^{k+i}+(k+i)y_{j}^{k+i}d_{dR}x_{j}^{-i}.$ And, the form given above can be obtained by replacing $\phi$ by $\phi+d_{dR}(\cdots).$ $\phi=k\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}x_{j}^{-i}d_{dR}y_{j}^{k+i}$ using the coordinates above. Combining these with the fact that $d_{dR}\phi=kd_{dR}\alpha$, we may represent $\alpha$ as
$\alpha/k=d_{dR}z^{k}+\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}x_{j}^{-i}d_{dR}y_{j}^{k+i},$ (3.26)
with its kernel generated by the vectors fields of the form $\partial/\partial x_{j}^{-i}\text{ and }\partial/\partial y^{k+i}_{j}-x^{-i}_{j}\partial/\partial z^{k}$. Here, the graded variables $x_{j}^{-i},y_{j}^{k+i},z^{k}$ on $\bf U$ serve as the desired local contact Darboux coordinates and we call $\alpha$ a $k$-contact element as in Example 3.10. Leaving $d_{dR}\alpha$ unchanged, we can modifying $H,\phi$ if necessary, so that we can also represent $\alpha$ as in Equation (3.18).
We then complete the proof of Theorem 3.13 when $k<0$ is odd.
When $k$ is not odd.
For the other cases $(a)\ k\equiv 0\mod 4,\text{ and }(b)\ k\equiv 2\mod 4$, one should use another sets of variables as in Equations ((a)) and ((b)), respectively, and modify $H,\phi,d$ as in Equations ((a)) -- ((b)). We leave details to the reader.
This completes the proof of Theorem 3.13 for all $k<0$, and hence that of Theorem 1.1.
Observation 3.17.
Denote by $B^{0}$ the subalgebra of $A^{0}$ with basis $x_{1}^{0},x_{2}^{0},\dots,x_{m_{0}}^{0},$ Then we define a sub-cdga $B$ of $A$ to be the free algebra over $B^{0}$ on generators $x_{j}^{-i},y_{j}^{k+i}$ only, with inclusion $\iota:B\hookrightarrow A.$ Observe that the elements $\phi,\omega^{0}:=d_{dR}\alpha|_{\ker\alpha}$ are all images under $\iota$ of the elements $\phi_{B},\omega_{B}^{0}:=\sum_{i=0}^{\ell}\sum_{j=1}^{m_{i}}d_{dR}x_{j}^{-i}d_{dR}y_{j}^{k+i}$, respectively. As in Section 2.3, $B$ is a minimal standard form cdga which in fact serves as a local symplectic model (for $k<0$ odd). As noted before, similar local models can be explicitly obtained for the other cases using Equations ((a)) -- ((b)).
In either case, suppose that we construct such $(A,B)$ for $k<0$, then the virtual dimension ${\rm vdim\ }B$ is always even, and hence the virtual dimension ${\rm vdim\ }A={\rm vdim\ }B-1$ is odd. In fact, if a cdga $A$ is in contact Darboux form, we have
${\rm vdim\ }A=\begin{cases*}-1+2\sum_{i}(-1)^{i}m_{i},&$k$ \mbox{\ even}\\ -1,&$k$ \mbox{ odd.}\end{cases*}$
4 Symplectifications
In this section, we give the formal description of the symplectification of a negatively shifted contact derived ${\mathbb{K}}$-scheme.
4.1 Symplectified spaces for contact manifolds
In classical contact geometry, for a contact manifold $(M,\xi=\ker(\alpha))$ with a globally defined contact 1-form $\alpha$, one can define the symplectification $\widetilde{M}$ of $M$ as the total space of the bundle $M\times\mathbb{R}^{*}\rightarrow M$ with a canonical symplectic form $\omega:=d_{dR}(e^{t}\alpha)$, where $t$ is the $\mathbb{R}$-coordinate. Most of the standard references [3, 17] use this approach, where the contact structure $\xi$ is in fact assumed to be coorientable (see Remark 3.2 for the definition).
However, for non-coorientible contact structures, the coordinate-dependent description above can no longer be applicable (as no global $\alpha$ and $t$-variable available); instead, we may use the following description from [2, Appendix 4.E]:  Given a contact manifold $(M,\xi)$, we let
$\widetilde{M}:=\big{\{}(p,\alpha_{p}):p\in M,\ \alpha_{p}\in T_{p}^{*}M,\ \text{s.t. }\ker\alpha_{p}=\xi_{p}\big{\}}.$ (4.1)
Here, $\widetilde{M}$ is just the set of all contact forms on the contact manifold. It should be noted that for a pair $(p,\alpha_{p})\in\widetilde{M}$, $\ \alpha_{p}$ is not a differential form but just a linear form on one tangent space $T_{p}M$ at the point of contact of the manifold such that its zero set is the contact plane. From [2, Appendix 4.E], it is straightforward to see that $\widetilde{M}$ is a smooth manifold of even dimension $\dim M+1$. Notice that there is a natural $\mathbb{R}^{*}$-action on $\widetilde{M}$ via $f\cdot(p,\alpha_{p})=(p,f\alpha_{p})$ such that $\widetilde{M}/\mathbb{R}^{*}\simeq M.$ Therefore, $\widetilde{M}$ can be identified as the total space of the $\mathbb{R}^{*}$-bundle over $M$.
From this identification with the $\mathbb{R}^{*}$-bundle $\pi:\widetilde{M}\rightarrow M$, the canonical symplectic structure on $\widetilde{M}$ can be defined as $\omega:=d_{dR}\lambda$, where the so-called canonical 1-form $\lambda$ is the differential 1-form on $\widetilde{M}$ whose value on any vector $v\in T_{x}\widetilde{M}$ at a point $x=(p,\alpha_{p})\in\widetilde{M}$ is given by
$\lambda_{x}(v):=\alpha_{p}(\pi_{*,x}(v)).$ (4.2)
4.2 Symplectified derived spaces
The construction of a canonical symplectified space associated to a contact space in terms of the total space of a line bundle with ${\mathbb{R}}^{*}$-action can be promoted to derived symplectic geometry. This leads to the definition of the symplectification of a shifted contact derived stack. In what follows, we explain the details. Let us begin by some relevant notions:
Definition 4.1.
Let ${\bf X}\in dStk_{{\mathbb{K}}}$ and $E\in QCoh(\bf X)$, then the total space $\bf\widetilde{E}$ of $E$ is defined as a derived stack sending
$A\mapsto\widetilde{{\bf E}}(A):=\{(p,s):\ p\in{\bf X}(A),\ s\in p^{*}E\},\text{ with }A\in cdga_{{\mathbb{K}}}.$ (4.3)
Here, from Yoneda’s lemma, ${\bf X}(A)\simeq Map_{dPstk}({\rm Spec}A,{\bf X})$, and hence any $A$-point $p\in{\bf X}(A)$ can be seen as a morphism $p:{\rm Spec}A\rightarrow\bf X$ of derived pre-stacks, and thus its pullback map $p^{*}:QCoh({\bf X})\rightarrow QCoh({\rm Spec}A)\simeq Mod_{A}$ sends $E\mapsto p^{*}E$. Hence, $s$ is just an element of the $A$-module $p^{*}E,$ a "fiber" over $p$.
Example 4.2.
if ${\bf X}\in dStk_{{\mathbb{K}}}$ admits a cotangent complex (which is always the case when $\bf X$ is also Artin), we can define the cotangent stack $\bf T^{*}X$ to be the total space $\bf\widetilde{{L}_{{\bf X}}}$ of $\mathbb{L}_{{\bf X}}\in QCoh(\bf X)$ and the $n$-shifted cotangent stack $\bf T^{*}[k]X$ to be the total space $\widetilde{\bf{L}_{{\bf X}}[k]}$ of $\mathbb{L}_{{\bf X}}[k]\in QCoh(\bf X)$. For more details see [15, §2].
Recall from [16] that for a perfect module $E$ over $\bf X$, its stack of sections $\widetilde{{\bf E}}$, defined by $\widetilde{{\bf E}}(-)=\mathbb{R}{\rm Spec}_{{\bf X}}(Sym(E^{\vee}))(-)$ is acted on by $\mathbb{G}_{m}$ because $Sym(E^{\vee})$ is graded $\mathcal{O}_{{\bf X}}$-algebra. This new grading is then called the fibre grading. Note also that the both zero section $\bf X\rightarrow\widetilde{E}$ and the projection $\bf\widetilde{E}\rightarrow X$ are $\mathbb{G}_{m}$-equivariant for the trivial $\mathbb{G}_{m}$-action on $\bf X.$ With this terminology, $\bf T^{*}[k]X$ is nothing but the stack of $k$-shifted 1-forms on $\bf X$ with a natural $\mathbb{G}_{m}$-action.
More generally, for $E\in QCoh(\bf X)$, the $\mathbb{G}_{m}$-action is given as a morphism of derived ${\mathbb{K}}$-stacks $\triangleleft:\mathbb{G}_{m}\times_{{\rm spec}{\mathbb{K}}}\bf\widetilde{E}\rightarrow\widetilde{E}$ such that for each $A\in cdga_{{\mathbb{K}}}$, we have a $\mathbb{G}_{m}(A)$-action $\triangleleft_{A}$ on ${\bf\widetilde{E}}(A)$, where $\mathbb{G}_{m}$ is the functor that maps $A\mapsto A^{\times}$. This definition also holds for any derived $S$-stack.
Now, we are in place of introducing the definition of the symplectification of a $k$-shifted contact derived stack using the machinery above:
Definition 4.3.
Let $\bf X$ be a locally finitely presented derived ${\mathbb{K}}$-scheme with a $k$-shifted contact structure $(\mathcal{K},\kappa,L)$. The symplectification is the total space $\widetilde{\bf L}$ of the $\mathbb{G}_{m}$-bundle of $L$, with a canonical $k$-shifted symplectic structure (for which the $\mathbb{G}_{m}$-action is of weight 1) as defined below.
4.2.1 Step-1: The derived stack $\mathcal{S}_{{\bf{X}}}$ of contact forms
Let $({\bf X};\mathcal{K},\kappa,L)$ be a $k$-shifted contact derived ${\mathbb{K}}$-scheme of locally finite presentation. Given $k<0$ and $p\in\bf X$, find a minimal standard form cdga $A$ and an affine derived sub-scheme ${\bf U}:={\rm Spec}A$ such that $p:{\rm Spec}A\rightarrow\bf X$ is Zariski open inclusion. Here, we assume w.l.o.g. that $L$ is trivial on $\bf U$.
Define a functor $\mathcal{S}_{{\bf{X}}}:cdga_{{\mathbb{K}}}\rightarrow Spcs$ by $A\mapsto\mathcal{S}_{{\bf{X}}}(A),$ where
$\mathcal{S}_{{\bf{X}}}(A):=\big{\{}(p,\alpha,v):p\in{\bf X}(A),\ \alpha:p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k],\ v:Cocone(\underline{\alpha})\xrightarrow{\sim}p^{*}(\mathcal{K})\big{\}},$ (4.4)
where each $v$ is a quasi-isomorphism respecting the natural morphisms $p^{*}\kappa:p^{*}\mathcal{K}\rightarrow p^{*}(\mathbb{T}_{{\bf X}})$ and $Cocone(\underline{\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})$. Under the current assumptions, the perfect complexes $\mathbb{T}_{A},\mathbb{L}_{A}$, when restricted to ${{\rm spec}H^{0}(A)}$, are both quasi-isomorphic to free complexes of $H^{0}(A)$-modules. For $A\in cdga_{{\mathbb{K}}}$, we then define a $\mathbb{G}_{m}(A)$-action on $\mathcal{S}_{{\bf{X}}}(A)$ by
$f\triangleleft(p,\alpha,v):=(p,f\cdot\alpha,v).$
Now, the following observation endows $\mathcal{S}_{{\bf{X}}}$ with the structure of a derived stack:
Proposition 4.4.
$\mathcal{S}_{{\bf{X}}}$ is equivalent to the total space $\bf\widetilde{L}$ of the $\mathbb{G}_{m}$-bundle of $L$ [^7] In general, when the group scheme $G=GL_{n}$, there is an equivalence between locally free sheaves of ${\rm rank\ }n$ and $GL_{n}$-torsors (hence principal $GL_{n}$-bundles). In this regard, a line bundle $L$ is nothing but a $\mathbb{G}_{m}$-bundle.. Therefore, it has the structure of a derived stack together with the projection map $\pi:\mathcal{S}_{{\bf{X}}}\rightarrow\bf X$. We then call $\mathcal{S}_{{\bf{X}}}$ the derived stack of $k$-contact forms.
Proof.
Let $(p,\alpha,v)$ be a point in $\mathcal{S}_{{\bf{X}}}(A)$. From definitions, we have two homotopy fiber sequences $(i)\ Cocone(\underline{\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})\rightarrow{\rm Im}\alpha$ and $(ii)\ Cocone({\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k]$. Then we get the following observation:
Observation 4.5.
There exists a triangle $Cocone(\underline{\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k]$.
Proof of Observation 4.5.
From the first sequence $(i)$ above, we have the following homotopy commutative diagram, where the left-hand square is the homotopy fiber.
[…]
There is a homotopy $H$ with $\underline{\alpha}\circ\pi_{1}\sim 0$. Using the monomorphism $j$, we get a homotopy $j\circ H$ such that $j\circ\underline{\alpha}\circ\pi_{1}\sim 0$, where $j\circ\underline{\alpha}\sim\alpha$. It follows that $Cocone(\underline{\alpha})$ homotopy commutes the outer diagram. Then the universality of $Cocone({\alpha})$ implies that $\exists\ \varphi_{1}:Cocone(\underline{\alpha})\rightarrow Cocone({\alpha})$.
Likewise, from the second sequence $(ii)$ above, we have the following homotopy commutative diagram.
[…]
Here, we have a homotopy such that $\alpha\circ\pi_{1}\sim 0.$ From the epi-mono factorization of $\alpha$, we have $j\circ\underline{\alpha}\circ\pi_{1}\sim 0$ as well. From $j\circ 0\sim 0$, we obtain $j\circ\underline{\alpha}\circ\pi_{1}\sim j\circ 0.$ Since $j$ is a monomorphism, we get the induced homotopy such that $\underline{\alpha}\circ\pi_{1}\sim 0.$ It follows from the universality of $Cocone(\underline{\alpha})$ that there exists a map $\varphi_{2}:Cocone({\alpha})\rightarrow Cocone(\underline{\alpha})$.
Using the maps $\varphi_{1},\varphi_{2}$ with the exact triangle $Cocone({\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k]$, we obtain a triangle $Cocone(\underline{\alpha})\rightarrow p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k]\rightarrow Cocone(\underline{\alpha})[1]$ as desired. ∎
Now, using Observation 4.5, we get an equivalence of triangles of perfect complexes on ${\rm Spec}A$
${p^{*}(\mathcal{K})}$${p^{*}(\mathbb{T}_{{\bf X}})}$${p^{*}(L)[k]}$${Cocone(\underline{\alpha})}$${p^{*}(\mathbb{T}_{{\bf X}})}$${\mathcal{O}[k],}$$\alpha$ $\simeq_{v}$ $id$ $\simeq$ (4.7)
and hence an induced isomorphism $p^{*}(L)\simeq\mathcal{O}.$ This map identifies $\mathcal{S}_{{\bf{X}}}$ with the total space of $L$.
4.2.2 Step-2: The induced shifted symplectic structure on $\mathcal{S}_{{\bf{X}}}$
Recall from [16] that a morphism $\bf Y\rightarrow\widetilde{E}$ of derived stacks, with $\bf\widetilde{E}$ is the total space of $E\in QCoh({\bf X})$, consists of a morphism $f:\bf Y\rightarrow X$ together with a section $s$ of $f^{*}E.$ If moreover, $\bf Y$ is a derived stack equipped with a $\mathbb{G}_{m}$-action, then a $\mathbb{G}_{m}$-equivariant morphism $\bf Y\rightarrow\widetilde{E}$ is given by the pair of a $\mathbb{G}_{m}$-equivariant morphism $f:\bf Y\rightarrow X$ and a $\mathbb{G}_{m}$-equivariant section $s$ of $f^{*}E\{1\}.$ In particular, the identity map $\bf\widetilde{E}\rightarrow\widetilde{E}$ corresponds to the pair of the projection map $\pi:\bf\widetilde{E}\rightarrow X$ and a ($\mathbb{G}_{m}$-equivariant) section of $\pi^{*}E\{1\}.$
Letting $\bf\widetilde{E}:=T^{*}[k]X$, the identity map $\bf T^{*}[k]X\rightarrow T^{*}[k]X$ is then determined by the data of the projection map $\pi_{{\bf X}}:\bf T^{*}[k]X\rightarrow X$ with a section of $\pi_{{\bf X}}^{*}\mathbb{L}_{{\bf X}}[k]$ (of weight 1 for the fiber grading). Since we have a natural map $\pi_{{\bf X}}^{*}\mathbb{L}_{{\bf X}}[k]\rightarrow\mathbb{L}_{{\bf T^{*}[k]X}}[k]$, this section induces a $k$-shifted 1-from $\lambda_{\bf X}$ on ${\bf T^{*}[k]X}$ called the tautological 1-form. Moreover, [16] shows that the induced closed 2-form $d_{dR}\lambda_{\bf X}$ of degree $k$ on ${\bf T^{*}[k]X}$ is in fact non-degenerate, and hence gives a $k$-shifted symplectic structure.
Definition 4.6.
By construction, we have the projection maps $\pi_{1}:\mathcal{S}_{{\bf{X}}}\rightarrow\bf X$ and $\pi_{2}:\mathcal{S}_{{\bf{X}}}\rightarrow{\bf T^{*}[k]X}.$ We define the canonical 1-from $\lambda$ on $\mathcal{S}_{{\bf X}}$ to be the pullback $\pi_{2}^{*}\lambda_{\bf X}$ of the tautological 1-form on ${\bf T^{*}[k]X}$.
Let $\mathcal{S}_{{\bf X}},\lambda$ be as above. Then ${\omega}:=(d_{dR}\lambda,0,0,\dots)$ is a $k$-shifted closed 2-form on $\mathcal{S}_{{\bf X}}$, and hence it defines a pre-$k$-shifted symplectic structure on $\mathcal{S}_{{\bf X}}$. Now, it remains to show that ${\omega}$ is non-degenerate.
Theorem 4.7.
Let $\bf X$ be a (locally finitely presented) derived ${\mathbb{K}}$-scheme with a $k$-shifted contact structure $(\mathcal{K},\kappa,L)$. The $k$-shifted closed 2-form ${\omega}$ described above is non-degenerate, and hence the derived stack $\mathcal{S}_{{\bf{X}}}\rightarrow\bf X$ is $k$-shifted symplectic.
We then call the pair $(\mathcal{S}_{\bf{X}},{\omega})$ the symplectification of $\bf X.$
Proof.
Given $k<0$ and $p\in\bf X$, apply Theorem 2.5 to get a refined neighborhood ${\bf U}={\rm Spec}A$ of $p$ with $q\in{\rm spec}H^{0}(A)$ such that $p:{\rm Spec}A\hookrightarrow\bf X$ is an open inclusion, with $q\mapsto p$, and $A$ is a minimal standard form cdga. W.l.o.g., we assume that $L$ is trivialized on $\bf U$, and hence, over $\bf U$, the induced 1-form $\alpha:p^{*}(\mathbb{T}_{{\bf X}})\rightarrow\mathcal{O}[k]$ is such that $Cocone(\underline{\alpha})\simeq p^{*}(\mathcal{K})$ and the 2-form $d_{dR}\alpha$ is non-degenerate on $p^{*}\mathcal{K}$. In that case, the fiber-cofiber sequence $p^{*}\mathcal{K}\rightarrow p^{*}\mathbb{T}_{{\bf X}}\rightarrow p^{*}(L[k])$ splits over $\bf U.$ We fix this locally defining 1-form $\alpha$ (and $u:Cocone(\underline{\alpha})\xrightarrow{\sim}p^{*}\mathcal{K}$) for the rest of the proof.
Denote the pullback of $\alpha$ under the open inclusion $p$ again by $\alpha\in p^{*}\mathbb{L}_{{\bf X}}[k]$. From definitions, the triple $(p,\alpha,u)$ is an element of $\mathcal{S}_{{\bf X}}(A)$.
Recall also that we have the distinguished triangle $p^{*}\mathbb{L}_{{\bf X}}\rightarrow\mathbb{L}_{A}\rightarrow\mathbb{L}_{p}$, where $\mathbb{L}_{p}$ is the relative cotangent complex, such that for refined neighborhoods, the restriction of $\mathbb{L}_{A}$ to ${\rm spec}H^{0}(A)$ is a free finite complex of $H^{0}(A)$-modules (cf. Prop. 2.15). Moreover, we have the identification $Cocone(\underline{\alpha})\simeq\ker\underline{\alpha}$, and $\ker\underline{\alpha}=\ker\alpha$ in $Mod_{A}$.
From Observations 3.7 & 3.9, both $\alpha$ and $f\cdot\alpha$ define equivalent contact structures, up to quasi-isomorphism, for any $f\in H^{0}(A)$ (after localizing $A$ at $q$ by $f$, if necessary).
It follows that, over ${\bf U}$, we can then identify the space $\mathcal{S}_{\bf{X}}$ locally as ${\bf U}\times_{{{\bf X}}}^{h}\mathbb{G}_{m},$ [^8] This also follows from the fact that $\mathcal{S}_{{\bf X}}$ is identified with the total space of $L$. Therefore, the corresponding homotopy fibers over $p$ are equivalent. with natural projections, where $\mathbb{G}_{m}={\rm Spec}B$ is the affine group scheme, with say $B:={\rm Spec}({\mathbb{K}}[x,x^{-1}])$. After localizing $A$ at $p$ if necessary, $\varphi\in{\mathbb{K}}[x,x^{-1}]$ acts on $f\in H^{0}(A)$ by $f\mapsto\varphi(f,f^{-1})$.
Recall that there exists a natural equivalence $DR(A)\otimes_{{\mathbb{K}}}DR(B)\simeq DR(A\otimes_{{\mathbb{K}}}B)$ induced by the identification
$\mathbb{L}_{A\otimes_{{\mathbb{K}}}B}\simeq(\mathbb{L}_{A}\otimes_{{\mathbb{K}}}B)\oplus(A\otimes_{{\mathbb{K}}}\mathbb{L}_{B}).$ (4.8)
Notice that for $q\in{\bf U}({\mathbb{K}})$, with $q:{\rm Spec}{\mathbb{K}}=\{\star\}\rightarrow\bf U$, and $f_{q}\in\mathbb{G}_{m}({\mathbb{K}})\simeq{\mathbb{K}}^{\times}$, $f_{q}\cdot\alpha$ is also a contact form at $q.$ Thus, on the part of the space $\mathcal{S}_{\bf{X}}$ over ${\bf U}$, we define a function $f$ with values in ${\mathbb{K}}^{\times}$. Then the canonical 1-form in Definition 4.6 can be locally written as
$\lambda=f\cdot\pi^{*}\alpha.$ (4.9)
Remark 4.8.
This local expression (4.9) follows from [18, Lemma 2.1.4] that the tautological 1-form on the shifted cotangent is universal in the sense that for any $k$-shifted 1-form $\beta$ on $\bf Y$ (viewed as $\beta:\bf Y\rightarrow T^{*}[k]Y$) we have $\beta^{*}\lambda_{{\bf Y}}=\beta$. Apply this universality for the case above ${\bf Y}:={\bf U}\times_{{{\bf X}}}^{h}\mathbb{G}_{m}$ with $\beta:=\pi^{*}\alpha$ and the $\mathbb{G}_{m}$-action to get the desired expression. We also have a non-zero factor $f$ as $\mathcal{S}_{{\bf X}}$ is identified with the total space of $L$ [^9] That is, for (non-zero) $\pi^{*}\alpha,\lambda\in\mathbb{L}_{p^{*}\mathcal{S}_{\bf{X}}}[k]$, this identification with the total space of $L$ (i.e. the space of trivializations) implies that there is a non-zero $f\in A$ s.t. $\lambda=f\cdot\pi^{*}\alpha.$
Lemma 4.9.
Locally on $\bf X$, for any locally defining 1-form $\alpha$, the $k$-shifted 2-form $\omega^{0}:=d_{dR}\lambda$ is non-degenerate, and hence the sequence $\omega:=(\omega^{0},0,0,\dots)$ defines a $k$-shifted symplectic structure on $\mathcal{S}_{\bf X}.$
Proof of Lemma 4.9.
Note first that the non-degeneracy in sense of Definition 2.19 can be equivalently formulated using refined local models as follows:
Observation 4.10.
A $k$-shifted 2-form $\gamma$ on ${\rm Spec}A$ for $A$ a minimal standard form cdga is non-degenerate if and only if for any non-zero vector $v\in\mathbb{T}_{A}|_{q\in{\rm spec}H^{0}(A)}$, there exists a non-zero vector $w\in\mathbb{T}_{A}|_{q\in{\rm spec}H^{0}(A)}$ such that $\iota_{w}\iota_{v}\gamma\neq 0.$ For a sketch, see the note  [^10] When restricted to ${\rm spec}H^{0}(A)$, the induced morphism $\mathbb{T}_{A}\rightarrow\Omega^{1}_{A}[k],\ Y\mapsto\iota_{Y}\gamma,$ in Definition 2.19 is just a map of finite complexes of free $H^{0}(A)$-modules. And, at $q\in{\rm spec}H^{0}(A)$, both $\mathbb{T}_{A}|_{q},\Omega^{1}_{A}|_{q}$ are complexes of ${\mathbb{K}}$-vector spaces. For non-degeneracy, we require the contraction map to be a (degree-wise) quasi-isomorphism. Recall that localizing $A$ at $q$ if necessary, we may assume that the induces map is indeed an (degree-wise) isomorphism near $q$. Therefore, Observation 4.10 is just an analogous result from linear algebra..
Now, to prove that $\omega^{0}:=d_{dR}\lambda$ is non-degenerate, we use Observation 4.10. That is, it suffices to show that, at $q\in\bf U,$ for any non-vanishing (homogeneous) vector field $\sigma\in(\mathbb{L}_{A\otimes_{{\mathbb{K}}}B})^{\vee}$, there is a vector field $\eta\in(\mathbb{L}_{A\otimes_{{\mathbb{K}}}B})^{\vee}$ such that $\iota_{\eta}(\iota_{\sigma}d_{dR}\lambda)\neq 0.$
For the rest of the proof, we will also assume that the shift $k$ is odd, say $k=-2\ell-1$ for $\ell\in{\mathbb{N}}$, to provide more explicit representations for vector fields of interest. In fact, our constructions will be independent of the choice of the shift and the corresponding graded variables.
Localizing $A$ at $q$ if necessary, choose the graded variables $x_{j}^{-i},y_{j}^{k+i},z^{k}$ on $\bf U$ as before so that $A$ is a standard form cdga over $A(0)$ freely generated by these graded variables, such that
$\ker\alpha|_{{\rm spec}H^{0}(A)}$ $=\big{\langle}\partial/\partial x_{1}^{-i},\dots,\partial/\partial x_{m_{i}}^{-i},\partial/\partial y^{k+i}_{1},\dots,\partial/\partial y^{k+i}_{m_{i}}:i=0,1,\dots,\ell\ \big{\rangle}_{H^{0}(A)},$
$Rest|_{{\rm spec}H^{0}(A)}$ $=\big{\langle}\partial/\partial z^{k}\big{\rangle}_{H^{0}(A)}.$
Observation 4.11.
As $f|_{q\in{\rm spec}H^{0}(A)}\in{\mathbb{K}}^{\times}$ and $\mathbb{L}_{A}=\Omega_{A}^{1}$ is a $A$-module, the first summand of Equation (4.8) can be equivalently written as, when restricted to ${{\rm spec}H^{0}(A)}$,
$\mathbb{L}_{A}\otimes_{{\mathbb{K}}}B|_{{\rm spec}H^{0}(A)}\simeq\langle d_{dR}x_{j}^{-i},d_{dR}y_{j}^{k+i},d_{dR}z^{k}\rangle_{H^{0}(A)}\otimes_{{\mathbb{K}}}{\mathbb{K}}[f,f^{-1}]\simeq\langle d_{dR}x_{j}^{-i},d_{dR}y_{j}^{k+i},d_{dR}z^{k}\rangle_{H^{0}(A)}.$
Using a cofibrant replacement of $B$ if necessary, the second summand of Equation (4.8) is just equivalent to $H^{0}(A)\otimes_{{\mathbb{K}}}\langle d_{dR}f\rangle_{B}.$
Now, using the natural splitting 3.21 and Observation 4.11 for the complex in Equation (4.8), we then have, when restricted to ${\rm spec}H^{0}(A)$,
$(\mathbb{L}_{A\otimes_{{\mathbb{K}}}B})^{\vee}\simeq\big{(}\ker\alpha\oplus Rest\big{)}\oplus\big{(}H^{0}(A)\otimes_{{\mathbb{K}}}\langle\partial/\partial f\rangle_{B}\big{)}.$ (4.10)
Using the splitting in Equation (4.10) with Observation 4.10, we prove the statement case by case. We first note that for any $\eta,\sigma\in(\mathbb{L}_{A\otimes_{{\mathbb{K}}}B})^{\vee}$ at $q$, direct computations give
$\iota_{\eta}(\iota_{\sigma}d_{dR}\lambda)=\mp(d_{dR}f)(\sigma)\alpha(\pi_{*}\eta)\mp(d_{dR}f)(\eta)\alpha(\pi_{*}\sigma)\mp f\cdot d_{dR}\alpha(\pi_{*}\sigma,\pi_{*}\eta).$
From Equation (4.10), it is enough to consider the following cases:
  1. 1.
    If $\sigma\in\ker\alpha$, then we have $\iota_{\eta}(\iota_{\sigma}d_{dR}\lambda)=\mp f\cdot d_{dR}\alpha(\sigma,\pi_{*}\eta)$. Since $d_{dR}\alpha|_{\ker\alpha}$ is non-degenerate by the contactness condition on $\alpha$ (and $f\in A^{\times}$), it is enough to take $\eta$ to be any non-zero vector in $\ker\alpha$.
  2. 2.
    If $\sigma\in Rest,$ then we get $\iota_{\eta}(\iota_{\sigma}d_{dR}\lambda)=\mp(d_{dR}f)(\eta)\alpha(\sigma).$ Here $\alpha(\sigma)\neq 0$ since $\sigma\in Rest.$ Thus, it is enough to take $\eta$ to be any non-zero vector in $H^{0}(A)\otimes_{{\mathbb{K}}}\langle\partial/\partial f\rangle_{B}$ so that $(d_{dR}f)(\eta)\neq 0.$
  3. 3.
    If $\sigma\in H^{0}(A)\otimes_{{\mathbb{K}}}\langle\partial/\partial f\rangle_{B},$ then $\iota_{\eta}(\iota_{\sigma}d_{dR}\lambda)=\mp(d_{dR}f)(\sigma)\alpha(\pi_{*}\eta).$ Observe that $(d_{dR}f)(\sigma)\neq 0,$ so it suffices to take $\eta$ to be any non-zero vector in $Rest$ so that $\alpha(\pi_{*}\eta)\neq 0.$
This completes the proof of Lemma 4.9, and hence that of Theorem 4.7.
Remark 4.12.
The proofs of Lemma 4.9 and Theorem 4.7 will still be valid for the other values of $k$. In fact, it clear to see that coordinates do not play any significant role in the proofs, rather than just providing explicit representations for the splitting.
In short, Lemma 4.9 has indeed a coordinate-free proof, and so does Theorem 4.7. Thus, using the same terminology as before, we say that the pair $(\mathcal{S}_{\bf{X}},\omega)$ above is the symplectification of the $k$-shifted contact derived ${\mathbb{K}}$-scheme $\bf X$ for any $k<0.$ Note also that this result is in fact canonical up to quasi-isomorphism by construction.
5 Concluding remarks
We conclude this paper with the following comments on the possible “stacky" generalizations of the main results presented in this paper and more.
Remark 5.1.
It should be noted that “stacky" generalizations of the results in [7] are also available in the literature. Ben-Bassat, Brav, Bussi and Joyce [9] extend the results of [7] from derived schemes to the case of (locally finitely presented) derived Artin ${\mathbb{K}}$-stacks. In short, Ben-Bassat, Brav, Bussi and Joyce [9] proved that derived Artin ${\mathbb{K}}$-stacks also have nice local models in some sense. Parts of the results from [9, Theorems 2.8 & 2.9] in fact give the generalization of Theorem 2.5 to the case of derived Artin ${\mathbb{K}}$-stacks. They also proved that every shifted symplectic derived Artin ${\mathbb{K}}$-stack admits the so-called “Darboux form atlas" [9, Theorem 2.10]. That is, their result extends Theorem 2.23 from derived ${\mathbb{K}}$-schemes to derived Artin ${\mathbb{K}}$-stacks.
Regarding the study of shifted contact structures, we note that Maglio, Tortorella and Vitagliano [19] have recently introduced and studied 00-shifted and $+1$-shifted contact structures on differentiable stacks, thus providing the foundations of shifted contact geometry in the stacky context.
In the sequels, our goals will be to extend the main results of this paper from derived schemes to the more general case of derived Artin stacks and to discuss more on the theory of shifted contact derived spaces, such as Legendrians and their local behavior, and sample constructions of contact stacks. In that respect, we propose:
Conjecture 5.2.
Theorem 3.13 and Theorem 4.7 still hold for (locally finitely presented) $k$-shifted contact derived Artin ${\mathbb{K}}$-stacks with $k<0$.
The foundational trilogy.
This is the first in a series of papers [20, 21]. There are other ongoing projects and more to come… In brief, the second paper [20] proves the conjecture above and gives several non-trivial constructions of contact derived stacks. In [21], we formally introduce Legendrians in the derived context and prove a Legendrian-Darboux-type theorem for the Legendrians in contact derived schemes.
Addendum.
In the latest version (Nov. 2024) of this document, Example 3.10 and the proof of Theorem 3.13 have been revisited (along with minor edits). The need for additional corrections and clarifications has emerged during the reviewing process of [20]. Similar modifications for [20] will appear in its final version. For details, see the addendum on the author’s webpage.
I wish to warmly thank the anonymous referee(s) for their valuable comments and suggestions, which helped a lot and improved the quality of the original manuscript(s).